### v1

# Differentially private and decentralized randomized power method

Abstract
The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks. As modern datasets contain sensitive private information, we need to give formal guarantees on the possible privacy leaks caused by this method. This paper focuses on enhancing privacy-preserving variants of the method. We propose a strategy to reduce the variance of the noise introduced to achieve Differential Privacy (DP). We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy. We leverage Secure Aggregation (a form of Multi-Party Computation) to allow the algorithm to perform computations using data distributed among multiple users or devices, without revealing individual data. We show that it is possible to use a noise scale in the decentralized setting that is similar to the one in the centralized setting. We improve upon existing convergence bounds for both the centralized and decentralized versions. The proposed method is especially relevant for decentralized applications such as distributed recommender systems, where privacy concerns are paramount.
Machine Learning, ICML
1 Introduction
The randomized power method has attracted significant attention due to its simplicity and its efficiency in solving large-scale linear algebra problems in spectral analysis or recommendation. It is in particular useful for computing the principal singular vectors of a matrix. It has a straightforward implementation which runs with low computational overhead (nearly linear complexity), making it particularly suitable for analyzing large modern datasets.
In large-scale machine learning systems, protecting user privacy is paramount. Datasets often contain sensitive personal or organizational information, so privacy has become a critical requirement. The standard randomized power method does not inherently provide privacy guarantees. While its output might seem less sensitive than the input data, there is no formal guarantee that private information embedded in the data cannot be inferred.
To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides formal guarantees and helps to mitigate potential privacy leaks of an algorithm. DP ensures that the output of an algorithm reveals little about any individual record in the input, even in the worst case. Several works have attempted to apply DP to the randomized power method. For instance, (Hardt & Price, 2014) and (Balcan et al., 2016) introduced differentially private versions of the method, and (Guo et al., 2021; Wang & Xu, 2020) explores federated approaches to private power iteration. Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
Despite these advancements, existing approaches suffer from several limitations. First, their performance heavily depends on the number of singular vectors being computed (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021; Liu et al., 2022; Wang & Xu, 2020), which impacts both utility and privacy guarantees. Second, they are primarily designed for centralized settings (Hardt & Price, 2014; Liu et al., 2022), where a trusted curator is assumed to hold the data. Moreover, some methods (Liu et al., 2022) make strong assumptions about the underlying data distribution (i.e. that the data is sub-Gaussian) which makes it harder to use these methods in practice. Some federated versions (Briguglio et al., 2023) claim to guarantee privacy by use of the federated setting, although it has been shown that decentralization does not offer privacy by design (Geiping et al., 2020). Finally, no fully decentralized versions exist to our knowledge. The previously proposed federated versions (Balcan et al., 2016; Guo et al., 2021; Wang & Xu, 2020) use public channel communication and ignore advances in Multi-Party-Computation (MPC). This makes them unsuitable for decentralized environments, such as distributed recommender systems and social networks, where data is partitioned across users and/or devices and communications are restricted to a predefined communication graph.
1.1 Contributions
The core focus of this article is to introduce improved privacy-preserving randomized power methods and to extend them to a decentralized setting, while maintaining computational efficiency and performance. We propose new convergence bounds for the (decentralized) privacy-preserving power method; these show that our proposed algorithm has a weaker dependence on the number of singular vectors that are computed.
We propose an improved differentially private version of the randomized power method, with tighter convergence bounds, reducing the noise required to achieve privacy and therefore improving the accuracy of the method. In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method. [^1] This proof addresses some minor mistakes in a privacy proof from (Hardt & Roth, 2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016). The modified result allows us to use a wider range of privacy parameters.
Finally, we propose decentralized versions of the algorithms. These are particularly relevant in contexts where data is distributed across multiple users or devices, as is often the case in recommender systems. To address the privacy concerns of the users, without forcing them to rely on a trusted curator, we propose a federated privacy-preserving version of the method using Secure Aggregation (Bell et al., 2020), a form of Multi-Party Computation (MPC) and a fully decentralized version using GOPA (GOssip noise for Private Averaging) (Sabater et al., 2022). This allows users to keep their data local while contributing to a privacy-preserving global computation. Our decentralized approach preserves the effectiveness of the centralized version of the randomized power method while incorporating the privacy advantages of local differential privacy (DP). Notably, by employing Secure Aggregation, we are able to set the noise scale comparable to that of central DP, enhancing accuracy. This makes the proposed method a promising candidate for decentralized and privacy-conscious applications.
Our contributions can be summarized as follows:
  1. 1.
    We introduce a novel privacy-preserving randomized power method that improves upon existing convergence bounds, with a reduced sensitivity to the number of singular vectors.
  2. 2.
    We extend the method to a decentralized setting, utilizing Secure Aggregation to ensure privacy in a distributed environment without the need for a trusted curator.
  3. 3.
    We provide new privacy proofs that correct errors in previous works, ensuring rigorous privacy guarantees.
  4. 4.
    We propose a new adjacency model to guarantee the differential privacy of our method, to open the door to new and more realistic use cases.
  5. 5.
    We present both runtime-dependent and runtime-independent bounds for the proposed methods, demonstrating their utility through theoretical analysis and empirical validation.
2 Background Material
2.1 Notation:
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a real-valued positive semi-definite matrix, and let $k$ denote the target rank, where typically $k\ll n$. We also define $\eta$ as a small positive scalar that represents the precision of the approximation.
The eigenvalue decomposition of $\bm{A}$ is given by:
$\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$ (1)
where $\bm{U}$ is the matrix of eigenvectors of $\bm{A}$, and $\bm{\Lambda}$ is a diagonal matrix containing the corresponding eigenvalues on its diagonal.
Let $\mathcal{N}(\mu,\sigma^{2})^{n\times p}$ be a ($n\times p$) random matrix of which each element is a random variable independent and identically distributed (i.i.d.) according to a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
For any matrix $\bm{X}$, let its element-wise maximum norm be $\|\bm{X}\|_{\max}=\max_{i,j}\|\bm{X}_{ij}\|$, where $\bm{X}_{ij}$ is the $(i,j)$-th entry of matrix $\bm{X}$. We denote by $\bm{X}_{j:}$ the $j$-th row of $\bm{X}$.
We also define two coherence measures for the matrix $\bm{A}$:
  • The $\mu_{0}$-coherence of $\bm{A}$ is the maximum absolute value of its eigenvector entries, defined as $\mu_{0}(\bm{A})=\|\bm{U}\|_{\max}=\max_{i,j}|\bm{U}_{ij}|$.
  • The $\mu_{1}$-coherence of $\bm{A}$ is the maximum row sum of the absolute values of its eigenvector entries, defined as $\mu_{1}(\bm{A})=\|\bm{U}\|_{\infty}=\max_{i}(\sum_{j}|\bm{U}_{ij}|)$.
Finally, we will also use the QR decomposition of matrix obtained using the Gram-Schmidt procedure. Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, the QR decomposition factorizes it as $\bm{X}=\bm{Q}\bm{R}$, where $\bm{Q}$ is an orthonormal matrix (i.e., $\bm{Q}^{\top}\bm{Q}=\bm{I}$) and $\bm{R}$ is an upper triangular matrix.
2.2 Differential privacy:
2.2.1 Approximate differential privacy:
A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (neighboring datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
$\Pr[\mathcal{M}(D_{1})\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(D_{2})\in S]+\delta,$ (2)
where $\epsilon$ is a small positive parameter representing the privacy loss (smaller values indicate stronger privacy guarantees), and $\delta$ is a small probability representing the chance that the privacy guarantee may fail, allowing for a small probability of privacy breach.
2.2.2 $\ell_{2}$-sensitivity:
The $\ell_{2}$-sensitivity of a function $f$, denoted as $\Delta_{2}(f)$ or $\Delta_{2}$ for brevity, measures the maximum change in the output of $f$ when its input changes by at most one element (for neighboring datasets $D_{1}\sim D_{2}$). It is defined as:
$\Delta_{2}=\max_{D_{1}\sim D_{2}}\|f(D_{1})-f(D_{2})\|_{2},$ (3)
where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm. This sensitivity quantifies the stability of the function under small changes in the input data and allows us to calibrate the noise needed to ensure differential privacy. The notion of adjacency of the datasets is in general application-dependent. Its granularity reflects what one wants to protect by using Differential Privacy. In (Hardt & Price, 2014; Balcan et al., 2016), it is defined as a magnitude 1 one-element change in the matrix of interest. In the context of recommender system, this would mean that we want to protect individual user-item interactions.
2.2.3 Zero-Concentrated Differential Privacy (zCDP):
A randomized algorithm $\mathcal{M}$ is said to satisfy $\rho$-zero-Concentrated Differential Privacy (zCDP) if for all neighboring datasets $D_{1}$ and $D_{2}$, and for all $\alpha\in(1,\infty)$, the following holds:
$D_{\alpha}(\mathcal{M}(D_{1})\|\mathcal{M}(D_{2}))\leq\rho\alpha\,,$ (4)
where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees). zCDP ensures that the privacy loss is tightly bounded in terms of the Rényi divergence, providing a more refined privacy analysis compared to pure or approximate differential privacy. Additionally, zCDP composes more easily than other differential privacy definitions and will be useful for our analysis.
The following lemma, introduced by Bun & Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
Lemma 2.1 (Gaussian Mechanism (Proposition 1.6 (Bun & Steinke, 2016))).
Let $f:X^{n}\to\mathbb{R}$ be a sensitivity-$\Delta$ function. Consider the mechanism (randomized algorithm) $M:X^{n}\to\mathbb{R}$, defined as $M(x)=f(x)+Z_{x}$ where for each input $x$, $Z_{x}$ is independently drawn from $\mathcal{N}(0,\sigma^{2})$. Then $M$ satisfies $\left(\frac{\Delta^{2}}{2\sigma^{2}}\right)$-zCDP.
The next lemma, which is a generalized version of a result presented by Bun & Steinke (2016), explains how a randomized algorithm, constructed by recursively composing a sequence of zCDP-satisying randomized algorithms, also satisfies zCDP.
Lemma 2.2 (Adaptive composition (Generalization from Lemma 2.3 (Bun & Steinke, 2016))).
Let $M_{1}:X^{n}\to Y_{1}$, $M_{2}:X^{n}\times Y_{1}\to Y_{2}$, …, $M_{L}:X^{n}\times Y_{1}\times\dots\times Y_{L-1}\to Y_{L}$ be randomized algorithms. Suppose $M_{i}$ satisfies $\rho_{i}$-zCDP as a function of its first argument for each $i=1,2,\dots,L$. Let $M^{\prime\prime}:X^{n}\to Y_{L}$, constructed recursively by:
$M^{\prime\prime}(x)=M_{L}(x,M_{L-1}(x,\dots,M_{2}(x,M_{1}(x))\dots)).$ (5)
Then $M^{\prime\prime}$ satisfies $(\sum_{i=1}^{L}\rho_{i})$-zCDP.
2.3 Privacy preserving power iteration:
Let $k$ be a target rank and $\eta$ a small real controlling the approximation precision. The aim of the randomized power method is to find a matrix $\bm{Q}\in\mathbb{R}^{n\times k}$ such that $\|\bm{U}-\bm{Q}\bm{Q}^{\top}\bm{U}\|\leq\eta$. In other words, we want $\bm{Q}$ to approximate the $k$ principal eigenvectors of $\bm{A}$. The aim of the privacy-preserving randomized power method is to compute $\bm{Q}$ without leaking sensitive information embedded in $\bm{A}$. In (Hardt & Price, 2014; Balcan et al., 2016), a sensitive change is modeled as a single-element change of maximum Frobenius norm 1 in $\bm{A}$. According to (Hardt & Price, 2014; Balcan et al., 2016), given approximate-DP parameters ($\epsilon,\delta$) we can compute $\bm{Q}$ using Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$.
1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$.
2:  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
3:  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\sigma=\epsilon^{-1}\sqrt{{4L\log(1/\delta)}}$;
4:  for $\ell=1$ to $L$ do
5:      1. Compute $\bm{Y}_{\ell}=\bm{A}\bm{X}^{\ell-1}+\bm{G}_{\ell}$ with $\bm{G}_{\ell}\sim\mathcal{N}(0,\sigma_{l}^{2}={\Delta}_{l}^{2}\cdot\sigma^{2})^{n\times p}$2. Compute QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
6:  end for
Algorithm 1 Privacy-preserving randomized power method
In recent works on the privacy-preserving power iteration (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021), the standard deviation of the noise related to Differential Privacy (DP) at iteration $l$ is:
$\sigma_{l}$ $=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$ (6)
$=\Delta_{l}^{a}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)},$ (7)
as defined in Fig. 3 of (Hardt & Price, 2014), Alg. 2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021). $\Delta_{l}^{a}$ is an $\ell_{2}$-sensitivity bound of the change in $\bm{A}\bm{X}^{\ell-1}$ given a change of magnitude at most 1 in $\bm{A}$ on only one element ($\Delta_{l}\leq\Delta_{l}^{a}$). Our first contribution comes from the fact that we can find a significantly tighter bound on $\Delta_{l}$.
2.4 Existing convergence bounds
To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in (Balcan et al., 2016). In Theorem 2.4, we present a reduction of their Theorem 3.1. The theorem, and several others we present relies on the verification of the conditions in Assumption 2.3.
Assumption 2.3.
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Let us fix:
$\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right).$ (8)
Assume that at every iteration $l$ of Algorithm 1, $\bm{G}_{\ell}$ satisfies, for some constant $\tau>0$:
$\|\bm{G}_{\ell}\|_{2}$ $=O\left(\eta(\sigma_{k}-\sigma_{q+1})\right)\,,\quad\text{and}$ (9)
$\quad\|\bm{U}_{q}^{\top}\bm{G}_{\ell}\|_{2}$ $=O\left(\eta\left(\sigma_{k}-\sigma_{q+1}\right)\frac{\sqrt{p}-\sqrt{q-1}}{\tau\sqrt{d}}\right)\,.$ (10)
Theorem 2.4 (Private Power Method (PPM) (Balcan et al., 2016)).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{l}{(\|\bm{X}_{l}\|_{\max})}\sqrt{pLd\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right)$
and also the following runtime-independent convergence bound
$\eta=O\left(\frac{\epsilon^{-1}\|\bm{U}\|_{\max}\sqrt{pLd\log(1/\delta)\log(d)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$
3 Analysis of Runtime-dependent Bounds
3.1 Proposed runtime-dependent bounds
3.1.1 Adjacency notion:
In (Hardt & Price, 2014; Balcan et al., 2016), a matrix $\bm{A}^{\prime}$ is denoted as adjacent to $\bm{A}$ ($\bm{A}\sim\bm{A}^{\prime}$) when $\bm{A}^{\prime}$ has a change of a single element of Frobenius norm at most $1$ compared to $\bm{A}$. In other words, we can write:
$\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top},$ (11)
with $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{d}$ being the canonical basis vector of dimension $d$. Without loss of generality, we have $0\leq i,j<n$ (the change of magnitude $c\leq 1$ can happen on any element $\bm{A}_{ij}$). In terms of differential privacy, this concept of adjacency allows us to protect element-wise updates to $\bm{A}$. We however note that as $\bm{A}$ and $\bm{A}^{\prime}$ are real and positive semi-definite, the change should also need to be on the diagonal of $\bm{A}$. We propose a new, more general, notion of adjacency to allow for other types of updates:
$\bm{A}^{\prime}=\bm{A}+\bm{C},$ (12)
with $\bm{C}$ a symmetric matrix representing the update, subject to $\sqrt{\sum_{i}\|C_{i:}\|_{1}^{2}}\leq 1$. The proposed adjacency notion is strictly more general than the previous adjacency notion (11). Indeed, $\bm{C}=c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top}$ is a special case. $\bm{C}$ can also for instance be non-zero on the diagonal or antidiagonal, or be another symmetric matrix.
3.1.2 Improved sensitivity bound:
The previous sensitivity bound $\Delta_{l}^{a}$ (used in (Balcan et al., 2016; Hardt & Price, 2014)) was found by using a straightforward extrapolation from the sensitivity of the change in $\bm{A}{\bm{X}^{\ell-1}}$ when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times 1}$ to when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$. We can however find a tighter bound on $\Delta_{l}$ if we directly analyse the change in ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$. We give the result of our improved analysis using our proposed adjacency notion (12) in Theorem 3.1 and provide its proof in Appendix A.1.
Theorem 3.1 (Improved Sensitivity Bound).
Let $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$ where $\bm{A}^{\prime}$ is defined using (12). Then $\Delta_{l}$ can be upped bounded as follows:
$\begin{split}\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{2}\triangleq\Delta_{l}^{b}\,.\end{split}$ (13)
3.1.3 Note on related privacy proofs:
Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by (Hardt & Price, 2014), (Balcan et al., 2016), (Hardt & Roth, 2013), (Hardt & Roth, 2012). These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP. Specifically, (Balcan et al., 2016) references the privacy proof from (Hardt & Price, 2014), which builds upon the privacy proof from (Hardt & Roth, 2013). The proof therein is also closely related to the one of (Hardt & Roth, 2012).
However, both (Hardt & Roth, 2013) (Theorem 2.4) and (Hardt & Roth, 2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped. This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all. The original composition rule is presented in Theorem III.3 of (Dwork et al., 2010). Moreover, Lemma 3.4 of (Hardt & Roth, 2013) misuses Theorem 2.4. Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration. By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
3.1.4 Privacy proof:
To address these issues, we propose a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2. This proof clarifies the ambiguities or errors present in the previous works, and ensures that our privacy guarantees are met.
Theorem 3.2 (Privacy proof for the PPM).
Assuming $\delta\in(0,1)$ and $\epsilon>0$ such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$, the Private Power Method with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ satisfies ($\epsilon,\delta$)-Differential Privacy.
3.1.5 Improved convergence bound:
Theorem 3.3 (Bound for the noisy power method (NPM) (Balcan et al., 2016)).
Let reals $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of a positive semi-definite matrix $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues and fix $\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right)$. If at every iteration $l$ of the NPM $\bm{G}_{\ell}$ satisfies Assumption 2.3, then after
$L=\Theta\left(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta.$ (14)
Theorem 3.4 (Improved PPM with runtime dependent bound).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$
Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (15)
Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016). Indeed the following bound holds:
[…]
which is tight only if $\forall j,\|\bm{X}_{ij}^{l}\|=\max_{j}\|\bm{X}_{:j}^{l}\|$. and can be in practice much tighter depending on $\bm{A}$. Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$. We provide the proof in Appendix A.2.
4 Analysis of Runtime-independent bounds
The runtime-independent bound that we develop in this section and present in Theorem 4.1 is a combination of two bounds. The first applies for matrices with large coherences, and is independent of the matrix coherence parameter. The second applies to matrices with low $\mu_{1}$-coherence, but the bound is dependent on the $\mu_{1}$-coherence of the matrix. We also show that our method achieves a bound which is at least as tight as both proposed bounds.
4.1 Proposed runtime-independent convergence bound
The proposed runtime-dependent bound that we presented earlier (Theorem 3.4) involves $\max_{i}\|\bm{X}_{i:}\|_{2}$, which depends on values that change during the execution of the algorithm. Similarly to (Hardt & Price, 2014; Hardt & Roth, 2013), we now propose a runtime-independent convergence bound for our method. This bound, presented in Theorem 4.1, is a combination of two bounds, one that is suited for matrices with high $\mu_{1}(\bm{A})$, and another that applies for matrices with small $\mu_{1}(\bm{A})$. We provide its proof in Appendix A.3. We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$. This is likely to happen in practice. We highlight this fact in Section 5.1.
Theorem 4.1.
Improved PPM with runtime independent bound
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$ Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 gives us the following runtime-independent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\cdot\min(1,\mu_{1}(\bm{A}))\cdot\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (17)
5 Empirical comparison of the proposed bounds
In previous sections, we proposed both runtime-dependent and runtime-independent bounds of convergence for our algorithm. The runtime-dependent bound is dependent on the matrix of interest $\bm{A}$ and on the matrices $\bm{X}_{l}$ that are computed iteratively. The runtime-independent bound (4.1) is dependent only on the matrix of interest $\bm{A}$. We now provide a few examples to justify the usefulness of the proposed bounds.
5.1 Empirical comparison of the runtime-independent bounds
As explained in Section 4.1, the proposed runtime-independent bound is tighter than those proposed in (Hardt & Price, 2014; Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$. We now provide an illustration of how the bounds compare for a practically relevant example, namely recommender systems.
Table 1: Statistics of datasets
Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{a}$ (existing) ${\Delta_{l}^{b}}$ (ours)
Gowalla 29,858 40,981 1,027,370 0.36 82.48 1.08 ×$\times\sqrt{p}$ 1
Yelp2018 31,668 38,048 1,561,406 0.40 60.42 1.29 ×$\times\sqrt{p}$ 1
Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.10 ×$\times\sqrt{p}$ 1
5.1.1 Recommender systems:
To compute several parameters of SOTA recommender system models, we need to compute leading eigenvectors of a matrix denoted as “the normalized item-item matrix”. This is equivalent as computing the right singular vectors of “the normalized adjacency matrix” (Shen et al., 2021). The most widely used datasets in recommender systems research are Gowalla (Cho et al., 2011), Yelp2018 (He et al., 2020) and Amazon-Book (Wang et al., 2019). Table 1 presents key statistics of these datasets, and compares the sensitivity bounds for the task of computing the leading eigenvectors of interest. As we can deduce from Table 1, the proposed bound allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
5.2 Empirical comparison of the runtime-dependent bounds
[…]
$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{a}]}{\mathbb{E}[\Delta_{0}^{b}]}.$ (18)
5.2.1 Theoretical approximation of $r(k,n)$:
We can find an asymptotic approximation of $r(k,n)$ using Approximation 5.1. The derivation is provided in Appendix A.4.
Approximation 5.1.
[…]
$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
5.2.2 Empirical approximation of $r(k,n)$:
Figure 1: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=8000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 5.1.
Figure 2: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=12000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 5.1.
We now want to compare theoretically and empirically how tight our proposed bound ($\Delta_{0}^{b}$) is compared to the previous one ($\Delta_{0}^{a}$). The scaling of the noise at the first iteration of the algorithm depends on the random initialization of the eigenvector matrix $\bm{X}_{0}$, which does not depend on the matrix of interest $\bm{A}$. We can therefore empirically estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$ on randomly initialized matrices $\bm{X}_{0}$ to compare the tightness of both bounds.
Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$. We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$. We use Theorem 5.1 to approximate it theoretically. To measure it empirically, we compute $\Delta_{0}^{a}$ and $\Delta_{0}^{b}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm. We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$. We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples. We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 5.1.
We give our results in Figures 1 and 2. We can see empirically and theoretically that the proposed noise scaling $\Delta_{l}^{b}$ is much tighter than $\Delta_{l}^{a}$ at the first step of the algorithm. Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice. The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
6 Decentralized version
Table 2: Overhead of using Secure Aggregation in our protocol
Client Server
Communication $O(log(s)+n^{2})$ $O(s\cdot log(s)+sn^{2})$
Computation $O(log^{2}(s)+n^{2}\cdot log(s))$ $O(s\cdot log^{2}(s)+sn^{2}\cdot log(s))$
In this section, we explore the scenario where the matrix $\bm{A}$ is distributed across multiple clients. Specifically, each client holds a matrix $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$ such that $\sum_{i=1}^{s}\bm{A}^{(s)}=\bm{A}$. The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(s)}$ to the server or other clients.
The randomized power method involves linear operations, which makes it well-suited for parallelization. An implementation of a private and distributed power method, using public channel communication between the clients and the server, was proposed in (Balcan et al., 2016). However, this approach relies on local Differential Privacy, since the data exchanged between the clients and the server can be observed by everyone.
To improve privacy while keeping the benefits of distributed computation, we propose using Secure Aggregation, a lightweight Multi-Party Computation (MPC) protocol. Secure Aggregation eliminates the need for local DP by allowing us to use distributed DP, which offers privacy guarantees similar to central DP. This approach has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
In distributed DP, each client adds noise to its own contributions before using the Secure Aggregation protocol. The noise is calibrated such that the sum of the noise across clients has a variance comparable to that used in Central DP, thus achieving similar privacy guarantees without requiring a trusted curator to aggregate the data.
We assume the “honest-but-curious” threat model, meaning that while the clients may attempt to learn additional information, they follow the protocol honestly. We also assume no corrupted or dropout users. For simplicity, this paper focuses on a basic version of distributed DP, neglecting effects of data quantization or errors introduced by modular arithmetic. We refer to (Kairouz et al., 2021) for an implementation of Distributed DP that takes these effects into account.
Finally, it is also possible to use a fully decentralized and differentially private protocol to perform averaging (Sabater et al., 2022). It is therefore straightforward to transform our protocol to avoid relying on a central server for Secure Aggregation, while keeping similar utility and communication costs.
6.1 Decentralized and private power method using distributed DP:
We give a federated version of the PPM in Algorithm 2. In this version, we reduce the variance of the noise by a multiplicative factor in O($s\cdot p\log({d})\cdot\|\bm{U}\|_{\infty}^{2}$) compared to (Balcan et al., 2016) due to the fact that we emulate Central DP with Secure Aggregation (thus using a secure channel) rather than using a public channel that only allows for local DP, and because we propose a generally tighter sensitivity bound. We show the equivalence between Algorithm 2 and Algorithm 1 in Theorem 6.2, and provide the proof in Appendix A.5. We first define the secure aggregation algorithm.
Definition 6.1.
We denote by $SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1{\leq}i{\leq}s\})$ the Secure Aggregation of locally held matrices $\bm{Y}_{\ell}^{(i)}$ over the set of users represented by indices $\{i|1{\leq}i{\leq}s\})$. It is equivalent to computing $\bm{Y}_{\ell}=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}$ over a secure channel.
  Input: distributed matrices $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, private parameters $\epsilon,\delta$.
  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\nu=\epsilon^{-1}\sqrt{\frac{4L\log(1/\delta)}{s}}$;
  for $\ell=1$ to $L$ do
      1. The central node broadcasts $\bm{X}^{\ell-1}$ to all $s$ computing nodes;2. Computing node $i$ computes $\bm{Y}^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ with $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$;3. The central node computes with the clients $\bm{Y}_{\ell}=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})$. 4. The central node computes QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
  end for
Algorithm 2 Federated private power method
Theorem 6.2 (Privacy and convergence of Algorithm 2).
Algorithm 2 is as privacy-preserving and has the same utility as its centralized version (Algorithm 1).
6.2 Communication and computation cost analysis:
We denote by $s$ the number of shards, such that $\sum_{i=1}^{s}\bm{A}^{(i)}=\bm{A}\in\mathbb{R}^{n\times n}$. Table 2 presents the communication and computations cost overhead introduced by the polylogarithmic version of Secure Aggregation that we employ (Bell et al., 2020).
7 Conclusion
We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems. We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees. By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, leading to higher utility without compromising on privacy. The decentralized variant of the algorithm maintains the efficiency and privacy of the centralized version but adapts it for distributed environments. Our theoretical and empirical results demonstrate the utility and privacy advantages of our approach, suggesting that it offers a strong solution for real-world decentralized applications.
8 Limitations
The proposed method is effective in ensuring differential privacy and decentralization but still has limitations. The privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic in certain environments. In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
Appendix A Proofs of Results
A.1 Proof of Theorem 3.1
Proof.
We note $\bm{A}^{\prime}$ a matrix adjacent to $\bm{A}$ using (12).
$\bm{A}^{\prime}=\bm{A}+\bm{C},$ (21)
with $\bm{C}$ a symmetric matrix representing the update, subject to $\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\leq 1$.
Then
$\begin{split}\Delta_{l}&=\|\bm{A}\bm{X}^{l}-\bm{A}^{\prime}\bm{X}^{l}\|_{F}\\ &=\|\bm{A}\bm{X}^{l}-(\bm{A}+\bm{C})\bm{X}^{l}\|_{F}\\ &=\|\bm{C}\bm{X}^{l}\|_{F}\\ &=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}.\\ \end{split}$ (22)
and
$\begin{split}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}&=\|\sum_{j}\bm{C}_{ij}\cdot\bm{e}_{j}^{\top}\cdot\bm{X}^{l}\|_{F}\\ &=\|\sum_{j}\bm{C}_{ij}\cdot\bm{X}^{l}_{j:}\|_{F}\\ &\leq\sum_{j}\|\bm{C}_{ij}\cdot\bm{X}^{l}_{j:}\|_{F}\\ &\leq\sum_{j}|\bm{C}_{ij}|\cdot\|\bm{X}^{l}_{j:}\|_{F}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|.\end{split}$ (23)
By injecting (22) into (23), we have:
$\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\Delta_{l}^{b}.\\ \end{split}$ (24)
Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\Delta_{l}^{b}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
A.2 Proof of Theorem 3.2
Proof.
We can see lines 4-5 of Algorithm 1 as a sequential composition ($M$) of $L$ Gaussian Mechanisms. By Lemma 2.1, each mechanism $M_{i}$ satisfies ($\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP where $\Delta_{i}$ is the $\ell_{2}$-sensitivity of the function associated to mechanism $\Delta_{i}$ and $\sigma_{i}^{2}$ is the variance of the noise added with the Gaussian Mechanism. By Lemma 2.2, the composition of mechanisms $M=(M_{1},...,M_{i},...,M_{L})$ satisfies ($\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP. Let $\rho\triangleq\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$. By design of our algorithm, we have:
$\rho$ $=\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$ (25)
$\leq\sum_{i=1}^{L}\frac{(\Delta_{l}^{b})^{2}}{2\sigma_{i}^{2}}$ (26)
$=\sum_{i=1}^{L}\frac{1}{2\sigma^{2}}$ (27)
$=\sum_{i=1}^{L}\frac{\epsilon^{2}}{{8L\log(1/\delta)}}$ (28)
$=\frac{\epsilon^{2}}{{8\log(1/\delta)}}.$ (29)
By Proposition 1.3 of (Bun & Steinke, 2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$. Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
Then:
$\epsilon^{\prime}$ $=\rho+2\sqrt{\rho\log(1/\delta)}$ (31)
$\leq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+2\sqrt{\frac{\epsilon^{2}}{{8\log(1/\delta)}}.\log(1/\delta)}$ (32)
$\leq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+\frac{\epsilon}{2}.$ (33)
To satisfy ($\epsilon$, $\delta$)-DP, we need:
$\epsilon\geq\epsilon^{\prime}$ $\impliedby\epsilon\geq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+\frac{\epsilon}{2}$ (35)
$\iff\frac{\epsilon}{2}\geq\frac{\epsilon^{2}}{{8\log(1/\delta)}}$ (36)
$\iff\epsilon\leq 4\log(1/\delta)$ (37)
$\iff\delta\leq\exp{(-\frac{\epsilon}{4})},$ (38)
which is a reasonable assumption, as in practice $\epsilon=O(1)$ and $\delta\ll\frac{1}{d}$ where $d$ is the number of records to protect. In our case, $d=n^{2}$ because we run the privacy-preserving power method on $\bm{A}\in\mathbb{R}^{n\times n}$. ∎
A.3 Proof of Theorem 3.4
Proof.
According to (Hardt & Price, 2014), if $\bm{G}_{l}\sim\mathcal{N}(0,\sigma_{l})^{d\times p}$, then with probability p=99/100 we have:
$\begin{split}&\max_{l=1}^{L}\|\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{d\cdot\log(L)}\,,\\ &\max_{l=1}^{L}\|\bm{U}^{\top}\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{p\cdot\log(L)}\,.\end{split}$ (40)
We can therefore satisfy the noise conditions of Theorem 3.3 with probability $p$ if we choose $\eta=\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}$, giving us:
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}\,.$ (41)
Which leads us to the following Theorem by injecting $\sigma_{l}=\Delta_{l}^{b}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
A.4 Derivation of Approximation 5.1
We provide here the derivation for Approximation 5.1:
Proof.
Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries. Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{i}sQ$ orthogonal and $\bm{R}$ upper triangular). By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$. Then by Theorem 2.2.1 of (Chikuse, 2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries. We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty. For a matrix $\bm{A}$, we note as $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation. Then we have:
$\bm{Q}_{ij}^{2}$ $=(\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})_{ij}^{2}$ (42)
$=(\bm{Z}^{\circ 2}((\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})^{\circ 2})_{ij}$ (43)
$=(\bm{Z}^{\circ 2}(\bm{Z}^{\top}\bm{Z})^{-1})_{ij}.$ (44)
We remark that each element of $\bm{Z}^{\circ 2}$ is distributed as a chi-squared variable with one degree of freedom, therefore $n\cdot\bm{Q}_{ij}^{2}$ is distributed as an i.i.d $F(1,n)$ random variable by definition. For large $n$, we can approximate these $F$ variables as Gaussians with mean $\mu=\frac{n}{n-2}$ and variances $\sigma^{2}=\frac{2n^{2}(n-1)}{(n-2)^{2}(n-4)}$. Therefore $\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{\mu}{n},\frac{\sigma^{2}}{n^{2}})$ and:
$k\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k^{2}\sigma^{2}}{n^{2}})$ $\sum_{j}\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k\sigma^{2}}{n^{2}}).$ (45)
We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 of (Massart, 1896). We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$. We assume that the variances of $\Delta_{l}^{a}$ and $\Delta_{l}^{b}$ (as $n$ and $k$ are large) are small and therefore:
$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\mathbb{E}[(\Delta_{l}^{a})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\mathbb{E}[(\Delta_{l}^{b})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
A.5 Proof of Theorem 4.1
We provide here the proof for Theorem 4.1: See 4.1
Proof.
First bound: We know by Section 2.4 of (Woodruff et al., 2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1. Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM. As proposed in (Hardt & Price, 2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar. Therefore we have:
$<x_{c},x_{o}>$ $=<\sum_{i=1}^{n}\alpha_{c}^{i}u_{i},\sum_{j=1}^{n}\alpha_{o}^{i}u_{j}>$ (48)
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}<u_{i},u_{j}>$ (By bilinearity) (49)
$=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}.$ (Orth. columns of $\bm{U}$) (50)
$\bm{X}$ has orthonormal columns, so if $c=o$, then $<x_{c},x_{o}>=\sum_{i=1}^{n}{(\alpha_{c}^{i})}^{2}=1$. Otherwise, we have $<x_{c},x_{o}>=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}=0$. We can therefore define a matrix $\bm{B}$ with orthonormal columns such that $\bm{X}=\bm{U}\bm{B}$ and $\bm{B}_{jc}=\alpha_{c}^{j}$. We can therefore write any row $i$ of $\bm{X}$ as $\bm{X}_{i:}=\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}$. Therefore
$\|\bm{X}_{i:}\|_{2}$ $=\|\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (51)
$\leq\sum_{j=1}^{p}\|\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (52)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|\|\bm{B}_{j:}\|_{2}$ (53)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|.$ (Orth. columns of $\bm{B}$) (54)
We then have $\max_{i}\|\bm{X}_{i:}\|_{2}\leq\max_{i}(\sum_{j=1}^{p}|\bm{U}_{ij}|)=\|\bm{U}\|_{\infty}=\mu_{1}(\bm{A})$.
We proposed two bounds for $\|\bm{X}_{i:}\|_{2}$, which are useful in two different cases. We can however note that:
$\|\bm{X}_{i:}\|_{2}\leq 1\text{ and }\|\bm{X}_{i:}\|_{2}\leq\mu_{1}(\bm{A})$ (55)
$\Rightarrow\|\bm{X}_{i:}\|_{2}\leq\min(1,\mu_{1}(\bm{A})),$ (56)
which leads us to Theorem 4.1 giving us a runtime-independent bound.
A.6 Proof of Theorem 6.2
Proof.
It is straightforward to see that if steps 2 and 3 of Algorithm 2 are equivalent to step 1 of Algorithm 1, then the two algorithms are equivalent. Recall that $Y^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ and $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$. Then steps 2 and 3 of Algorithm 2 correspond to:
$\begin{split}\bm{Y}_{\ell}&=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})\\ &=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}\\ &=\sum_{i=1}^{s}(\bm{A}^{(i)}\bm{X}_{\ell-1}+\bm{G}^{(i)}_{\ell})\\ &=\bm{A}\bm{X}^{\ell-1}+\sum_{i=1}^{s}\bm{G}^{(i)}_{\ell}\\ &=\bm{A}\bm{X}^{\ell-1}+\bm{G}_{\ell},\end{split}$ (57)
where $\bm{G}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\cdot(s\nu^{2}))^{n\times p}$, and we have $s\nu^{2}=\sigma^{2}$ by definition, completing the equivalence proof. ∎

### v2

@@
 The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks.
-As modern datasets contain sensitive private information, we need to give formal guarantees on the possible privacy leaks caused by this method.
-This paper focuses on enhancing privacy-preserving variants of the method.
-We propose a strategy to reduce the variance of the noise introduced to achieve Differential Privacy (DP).
+However, its application to large datasets containing personal information (e.g.
+web interactions, search history, personal tastes) raises critical privacy concerns.
+This paper addresses these issues by proposing enhanced privacy-preserving variants of the method, which reduce the variance of the noise introduced to achieve Differential Privacy (DP).
 We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy.
-We leverage Secure Aggregation (a form of Multi-Party Computation) to allow the algorithm to perform computations using data distributed among multiple users or devices, without revealing individual data.
-We show that it is possible to use a noise scale in the decentralized setting that is similar to the one in the centralized setting.
+We leverage Secure Aggregation to run the algorithm on data distributed among multiple users or devices, while using a noise scale in the decentralized setting that is similar to the one in the centralized setting yet without revealing individual data.
 We improve upon existing convergence bounds for both the centralized and decentralized versions.
-The proposed method is especially relevant for decentralized applications such as distributed recommender systems, where privacy concerns are paramount.
 Machine Learning, ICML
 1 Introduction
-The randomized power method has attracted significant attention due to its simplicity and its efficiency in solving large-scale linear algebra problems in spectral analysis or recommendation.
-It is in particular useful for computing the principal singular vectors of a matrix.
-It has a straightforward implementation which runs with low computational overhead (nearly linear complexity), making it particularly suitable for analyzing large modern datasets.
+The randomized power method is a simple and computationally efficient tool for solving large-scale linear algebra problems, particularly in spectral analysis and recommendation systems.
+It approximates the principal singular vectors of a matrix with low computational complexity (nearly linear), making it particularly suitable for analyzing large modern datasets.
 In large-scale machine learning systems, protecting user privacy is paramount.
@@
 While its output might seem less sensitive than the input data, there is no formal guarantee that private information embedded in the data cannot be inferred.
-To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides formal guarantees and helps to mitigate potential privacy leaks of an algorithm.
-DP ensures that the output of an algorithm reveals little about any individual record in the input, even in the worst case.
+To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides strong formal guarantees and mitigate potential privacy leaks of an algorithm.
+DP ensures that the output of an algorithm reveals little about any individual record in the input.
 Several works have attempted to apply DP to the randomized power method.
-For instance, (Hardt & Price, 2014) and (Balcan et al., 2016) introduced differentially private versions of the method, and (Guo et al., 2021; Wang & Xu, 2020) explores federated approaches to private power iteration.
+For instance, Hardt & Price (2014) and Balcan et al.
+(2016) introduced differentially private versions of the method, and Guo et al.
+(2021); Wang & Xu (2020) explored federated approaches to private power iteration.
 Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
@@
 In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method.
-[^1] This proof addresses some minor mistakes in a privacy proof from (Hardt & Roth, 2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016).
+[^1] This proof addresses some minor mistakes in a privacy proof from Hardt & Roth (2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016).
 The modified result allows us to use a wider range of privacy parameters.
@@
 This makes the proposed method a promising candidate for decentralized and privacy-conscious applications.
-Our contributions can be summarized as follows:
-  1. 1.
-    We introduce a novel privacy-preserving randomized power method that improves upon existing convergence bounds, with a reduced sensitivity to the number of singular vectors.
-  2. 2.
-    We extend the method to a decentralized setting, utilizing Secure Aggregation to ensure privacy in a distributed environment without the need for a trusted curator.
-  3. 3.
-    We provide new privacy proofs that correct errors in previous works, ensuring rigorous privacy guarantees.
-  4. 4.
-    We propose a new adjacency model to guarantee the differential privacy of our method, to open the door to new and more realistic use cases.
-  5. 5.
-    We present both runtime-dependent and runtime-independent bounds for the proposed methods, demonstrating their utility through theoretical analysis and empirical validation.
 2 Background Material
@@
 2.2.1 Approximate differential privacy:
-A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (neighboring datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
+A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (adjacent datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
 $\Pr[\mathcal{M}(D_{1})\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(D_{2})\in S]+\delta,$ (2)
@@
 where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm.
-This sensitivity quantifies the stability of the function under small changes in the input data and allows us to calibrate the noise needed to ensure differential privacy.
-The notion of adjacency of the datasets is in general application-dependent.
-Its granularity reflects what one wants to protect by using Differential Privacy.
-In (Hardt & Price, 2014; Balcan et al., 2016), it is defined as a magnitude 1 one-element change in the matrix of interest.
-In the context of recommender system, this would mean that we want to protect individual user-item interactions.
 2.2.3 Zero-Concentrated Differential Privacy (zCDP):
@@
 where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees).
-zCDP ensures that the privacy loss is tightly bounded in terms of the Rényi divergence, providing a more refined privacy analysis compared to pure or approximate differential privacy.
-Additionally, zCDP composes more easily than other differential privacy definitions and will be useful for our analysis.
 The following lemma, introduced by Bun & Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
@@
 2.3 Privacy preserving power iteration:
-Let $k$ be a target rank and $\eta$ a small real controlling the approximation precision.
-The aim of the randomized power method is to find a matrix $\bm{Q}\in\mathbb{R}^{n\times k}$ such that $\|\bm{U}-\bm{Q}\bm{Q}^{\top}\bm{U}\|\leq\eta$.
-In other words, we want $\bm{Q}$ to approximate the $k$ principal eigenvectors of $\bm{A}$.
-The aim of the privacy-preserving randomized power method is to compute $\bm{Q}$ without leaking sensitive information embedded in $\bm{A}$.
-In (Hardt & Price, 2014; Balcan et al., 2016), a sensitive change is modeled as a single-element change of maximum Frobenius norm 1 in $\bm{A}$.
-According to (Hardt & Price, 2014; Balcan et al., 2016), given approximate-DP parameters ($\epsilon,\delta$) we can compute $\bm{Q}$ using Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$.
+Let $k$ denote the target rank, $q$ a small real and $\eta$ a small positive value controlling the precision of the approximation.
+The goal of the randomized power method is to compute a matrix $\bm{Q}\in\mathbb{R}^{n\times p}$, where $p=k+q$, such that the range of $\bm{Q}$ approximates the range of the top $k$ eigenvectors of $\bm{A}$, which are denoted $\bm{U}_{k}$.
+Specifically, we aim to ensure that:
+$\|\bm{U}_{k}-\bm{Q}\bm{Q}^{\top}\bm{U}_{k}\|\leq\eta.$
+The privacy-preserving randomized power method extends this approach to protect sensitive information in $\bm{A}$ during computation.
+Prior work (Hardt & Price, 2014; Balcan et al., 2016) show that under approximate Differential Privacy (DP) parameters $(\epsilon,\delta)$, $\bm{Q}$ can be computed using Algorithm 1, where the sensitivity $\Delta_{l}$ at iteration $l$ is bounded by $\Delta_{l}^{\text{prior}}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$.
+Here, $\Delta_{l}^{\text{prior}}$ represents an $\ell_{2}$-sensitivity bound that quantifies the impact of a single-element change in $\bm{A}$ on $\bm{A}\bm{X}^{l-1}$.
+Specifically, $\Delta_{l}\leq\Delta_{l}^{\text{prior}}$.
+With this bound [^2] As defined in Fig.
+3 of (Hardt & Price, 2014), Alg.
+2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021)., the standard deviation of the noise $\sigma_{l}$ associated to DP at iteration $l$ is given by:
+$\sigma_{l}$ $=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$ (6)
+$=\Delta_{l}^{\text{prior}}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}.$ (7)
+As a first contribution, we refine this sensitivity analysis by introducing a significantly tighter bound for $\Delta_{l}$, reducing the noise required for DP and therefore allowing tighter convergence bounds (smaller $\eta$).
 1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$.
@@
 Algorithm 1 Privacy-preserving randomized power method
-In recent works on the privacy-preserving power iteration (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021), the standard deviation of the noise related to Differential Privacy (DP) at iteration $l$ is:
-$\sigma_{l}$ $=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$ (6)
-$=\Delta_{l}^{a}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)},$ (7)
-as defined in Fig.
-3 of (Hardt & Price, 2014), Alg.
-2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021).
-$\Delta_{l}^{a}$ is an $\ell_{2}$-sensitivity bound of the change in $\bm{A}\bm{X}^{\ell-1}$ given a change of magnitude at most 1 in $\bm{A}$ on only one element ($\Delta_{l}\leq\Delta_{l}^{a}$).
-Our first contribution comes from the fact that we can find a significantly tighter bound on $\Delta_{l}$.
 2.4 Existing convergence bounds
-To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in (Balcan et al., 2016).
+To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in Balcan et al.
+(2016).
 In Theorem 2.4, we present a reduction of their Theorem 3.1.
@@
 Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues.
-Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}$ gives us the following runtime-dependent convergence bound:
+Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}$ gives us the following runtime-dependent convergence bound:
 $\eta=O\left(\frac{\epsilon^{-1}\max_{l}{(\|\bm{X}_{l}\|_{\max})}\sqrt{pLd\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right)$
@@
 3.1.1 Adjacency notion:
-In (Hardt & Price, 2014; Balcan et al., 2016), a matrix $\bm{A}^{\prime}$ is denoted as adjacent to $\bm{A}$ ($\bm{A}\sim\bm{A}^{\prime}$) when $\bm{A}^{\prime}$ has a change of a single element of Frobenius norm at most $1$ compared to $\bm{A}$.
-In other words, we can write:
-$\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top},$ (11)
-with $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{d}$ being the canonical basis vector of dimension $d$.
-Without loss of generality, we have $0\leq i,j<n$ (the change of magnitude $c\leq 1$ can happen on any element $\bm{A}_{ij}$).
-In terms of differential privacy, this concept of adjacency allows us to protect element-wise updates to $\bm{A}$.
-We however note that as $\bm{A}$ and $\bm{A}^{\prime}$ are real and positive semi-definite, the change should also need to be on the diagonal of $\bm{A}$.
+In prior works (Hardt & Price, 2014; Balcan et al., 2016), two matrices $\bm{A}$ and $\bm{A}^{\prime}$ are considered adjacent (denoted $\bm{A}\sim\bm{A}^{\prime}$) if they differ by a change in a single element, with a Frobenius norm of at most 1:
+$\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\bm{e}_{j}^{\top},$ (11)
+where $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{n}$ are canonical basis vectors of dimension $d$, $c\leq 1$ represents the magnitude of the change, and $0\leq i,j<n$.
+This adjacency notion models a sensitive change as a modification to a single element in $\bm{A}$ to protect individual element-wise updates under Differential Privacy.
+However, since $\bm{A}$ is symmetric positive semi-definite and changes must preserve this property, this necessitates updates on the diagonal elements.
 We propose a new, more general, notion of adjacency to allow for other types of updates:
 $\bm{A}^{\prime}=\bm{A}+\bm{C},$ (12)
-with $\bm{C}$ a symmetric matrix representing the update, subject to $\sqrt{\sum_{i}\|C_{i:}\|_{1}^{2}}\leq 1$.
-The proposed adjacency notion is strictly more general than the previous adjacency notion (11).
-Indeed, $\bm{C}=c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top}$ is a special case.
-$\bm{C}$ can also for instance be non-zero on the diagonal or antidiagonal, or be another symmetric matrix.
+where $\bm{C}\in\mathbb{R}^{n\times n}$ is a symmetric matrix representing the update, subject to the condition:
+$\sqrt{\sum_{i=1}^{n}\|\bm{C}_{i,:}\|_{1}^{2}}\leq 1.$ (13)
+The proposed adjacency notion is strictly more general than Equation (11).
+For example, setting $\bm{C}=c\cdot\bm{e}_{i}\bm{e}_{j}^{\top}$ recovers the original definition.
+Additionally, $\bm{C}$ can have non-zero entries on the diagonal, anti-diagonal, or any symmetric pattern, allowing for a variety of updates while maintaining symmetry.
+In the context of recommender systems where $\bm{A}=\bm{R}^{\top}\bm{R}$ and $\bm{R}$ is a user-item interaction matrix, this adjacency notion allows for element-wise changes in $\bm{R}$ (i.e.
+protect user-item interactions).
+By adopting this more general adjacency definition, we make our privacy guarantees applicable to a wider range of real-world scenarios.
 3.1.2 Improved sensitivity bound:
-The previous sensitivity bound $\Delta_{l}^{a}$ (used in (Balcan et al., 2016; Hardt & Price, 2014)) was found by using a straightforward extrapolation from the sensitivity of the change in $\bm{A}{\bm{X}^{\ell-1}}$ when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times 1}$ to when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$.
-We can however find a tighter bound on $\Delta_{l}$ if we directly analyse the change in ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$.
-We give the result of our improved analysis using our proposed adjacency notion (12) in Theorem 3.1 and provide its proof in Appendix A.1.
+The previous sensitivity bound $\Delta_{l}^{\text{prior}}$ (Hardt & Price, 2014; Balcan et al., 2016) for $\bm{A}\bm{X}^{l-1}$ was estimated by extrapolating from the case where $\bm{X}^{l-1}\in\mathbb{R}^{n\times 1}$ to the general case $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$.
+This leads to an overestimation of the sensitivity and to unnecessarily large noise addition.
+We directly analyze the change in $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$ and derive a tighter bound on the sensitivity $\Delta_{l}$.
+By using our proposed adjacency notion from Equation (12), we establish the following result (proof provided in Appendix A.1):
 Theorem 3.1 (Improved Sensitivity Bound).
-Let $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$ where $\bm{A}^{\prime}$ is defined using (12). Then $\Delta_{l}$ can be upped bounded as follows:
-$\begin{split}\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{2}\triangleq\Delta_{l}^{b}\,.\end{split}$ (13)
-3.1.3 Note on related privacy proofs:
-Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by (Hardt & Price, 2014), (Balcan et al., 2016), (Hardt & Roth, 2013), (Hardt & Roth, 2012).
-These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP.
-Specifically, (Balcan et al., 2016) references the privacy proof from (Hardt & Price, 2014), which builds upon the privacy proof from (Hardt & Roth, 2013).
-The proof therein is also closely related to the one of (Hardt & Roth, 2012).
-However, both (Hardt & Roth, 2013) (Theorem 2.4) and (Hardt & Roth, 2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped.
-This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all.
-The original composition rule is presented in Theorem III.3 of (Dwork et al., 2010).
-Moreover, Lemma 3.4 of (Hardt & Roth, 2013) misuses Theorem 2.4.
-Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration.
-By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
+Let $\bm{A}^{\prime}$ be defined as in Equation (12), and consider the sensitivity $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$. Then, $\Delta_{l}$ is upped-bounded as follows:
+$\Delta_{l}\leq\max_{i}\|\bm{X}_{i,:}^{l}\|_{2}\triangleq\hat{\Delta}_{l}.$ (14)
+This improved bound $\hat{\Delta}_{l}$ is generally tighter than the previous bound $\Delta_{l}^{\text{prior}}=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$, as it eliminates the dependence on $\sqrt{p}$ and directly relates to the maximum row norms of $\bm{X}^{l}$.
+By reducing the estimated sensitivity, we can add less noise while still satisfying Differential Privacy, thereby improving the utility of the algorithm.
 3.1.3 Privacy proof:
-To address these issues, we propose a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2.
-This proof clarifies the ambiguities or errors present in the previous works, and ensures that our privacy guarantees are met.
+To address these issues, we propose a result with a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2.2.
+This proof clarifies the ambiguities or errors present in the previous works (see Appendix A.2), and ensures that our privacy guarantees are met.
 Theorem 3.2 (Privacy proof for the PPM).
@@
 $L=\Theta\left(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
-iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have
+iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have:
 $\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta.$ (15)
@@
 $\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (16)
-Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016). Indeed the following bound holds:
-[…]
-which is tight only if $\forall j,\|\bm{X}_{ij}^{l}\|=\max_{j}\|\bm{X}_{:j}^{l}\|$.
-and can be in practice much tighter depending on $\bm{A}$.
+Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016), and can be in practice much tighter depending on $\bm{A}$.
 Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$.
-We provide the proof in Appendix A.2.
+We provide the proof in Appendix A.2.2.
 4 Analysis of Runtime-independent bounds
-The runtime-independent bound that we develop in this section and present in Theorem 4.1 is a combination of two bounds.
-The first applies for matrices with large coherences, and is independent of the matrix coherence parameter.
-The second applies to matrices with low $\mu_{1}$-coherence, but the bound is dependent on the $\mu_{1}$-coherence of the matrix.
-We also show that our method achieves a bound which is at least as tight as both proposed bounds.
 4.1 Proposed runtime-independent convergence bound
-The proposed runtime-dependent bound that we presented earlier (Theorem 3.4) involves $\max_{i}\|\bm{X}_{i:}\|_{2}$, which depends on values that change during the execution of the algorithm.
-Similarly to (Hardt & Price, 2014; Hardt & Roth, 2013), we now propose a runtime-independent convergence bound for our method.
-This bound, presented in Theorem 4.1, is a combination of two bounds, one that is suited for matrices with high $\mu_{1}(\bm{A})$, and another that applies for matrices with small $\mu_{1}(\bm{A})$.
-We provide its proof in Appendix A.3.
+We presented in Theorem 3.4 a runtime-dependent convergence bound involving $\max_{i}\|\bm{X}_{i,:}\|_{2}$, which is only known during the execution of the algorithm.
+To provide a more practical analysis, we now derive a runtime-independent convergence bound in Theorem 4.1.
+This bound is a combination of two cases:
+  • High Coherence Case: Applies to matrices with large coherence parameters, and the bound is independent of the matrix’s coherence.
+  • Low Coherence Case: Applies to matrices with low $\mu_{1}$-coherence, with the bound depending explicitly on $\mu_{1}(\bm{A})$.
+We demonstrate that our method achieves a convergence bound that is at least as tight as both of these bounds.
 We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$.
-This is likely to happen in practice.
-We highlight this fact in Section 5.1.
+This is likely to happen in practice, as we highlight in Section 6.1.1.
 Theorem 4.1.
@@
 $\eta=O\left(\frac{\epsilon^{-1}\cdot\min(1,\mu_{1}(\bm{A}))\cdot\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (17)
-5 Empirical comparison of the proposed bounds
-In previous sections, we proposed both runtime-dependent and runtime-independent bounds of convergence for our algorithm.
-The runtime-dependent bound is dependent on the matrix of interest $\bm{A}$ and on the matrices $\bm{X}_{l}$ that are computed iteratively.
-The runtime-independent bound (4.1) is dependent only on the matrix of interest $\bm{A}$.
-We now provide a few examples to justify the usefulness of the proposed bounds.
-5.1 Empirical comparison of the runtime-independent bounds
-As explained in Section 4.1, the proposed runtime-independent bound is tighter than those proposed in (Hardt & Price, 2014; Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$.
-We now provide an illustration of how the bounds compare for a practically relevant example, namely recommender systems.
-Table 1: Statistics of datasets
-Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{a}$ (existing) ${\Delta_{l}^{b}}$ (ours)
-Gowalla 29,858 40,981 1,027,370 0.36 82.48 1.08 ×$\times\sqrt{p}$ 1
-Yelp2018 31,668 38,048 1,561,406 0.40 60.42 1.29 ×$\times\sqrt{p}$ 1
-Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.10 ×$\times\sqrt{p}$ 1
-5.1.1 Recommender systems:
-To compute several parameters of SOTA recommender system models, we need to compute leading eigenvectors of a matrix denoted as “the normalized item-item matrix”.
-This is equivalent as computing the right singular vectors of “the normalized adjacency matrix” (Shen et al., 2021).
-The most widely used datasets in recommender systems research are Gowalla (Cho et al., 2011), Yelp2018 (He et al., 2020) and Amazon-Book (Wang et al., 2019).
-Table 1 presents key statistics of these datasets, and compares the sensitivity bounds for the task of computing the leading eigenvectors of interest.
-As we can deduce from Table 1, the proposed bound allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
-5.2 Empirical comparison of the runtime-dependent bounds
-[…]
-$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{a}]}{\mathbb{E}[\Delta_{0}^{b}]}.$ (18)
-5.2.1 Theoretical approximation of $r(k,n)$:
-We can find an asymptotic approximation of $r(k,n)$ using Approximation 5.1. The derivation is provided in Appendix A.4.
-Approximation 5.1.
-[…]
-$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
-$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
-5.2.2 Empirical approximation of $r(k,n)$:
-Figure 1: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=8000$ and $k$ ranging between 64 and 4000, with a step of 64.
-Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm.
-The theoretical ratios are estimated using Theorem 5.1.
-Figure 2: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=12000$ and $k$ ranging between 64 and 4000, with a step of 64.
-Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm.
-The theoretical ratios are estimated using Theorem 5.1.
-We now want to compare theoretically and empirically how tight our proposed bound ($\Delta_{0}^{b}$) is compared to the previous one ($\Delta_{0}^{a}$).
-The scaling of the noise at the first iteration of the algorithm depends on the random initialization of the eigenvector matrix $\bm{X}_{0}$, which does not depend on the matrix of interest $\bm{A}$.
-We can therefore empirically estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$ on randomly initialized matrices $\bm{X}_{0}$ to compare the tightness of both bounds.
-Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$.
-We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$.
-We use Theorem 5.1 to approximate it theoretically.
-To measure it empirically, we compute $\Delta_{0}^{a}$ and $\Delta_{0}^{b}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm.
-We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$.
-We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples.
-We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 5.1.
-We give our results in Figures 1 and 2.
-We can see empirically and theoretically that the proposed noise scaling $\Delta_{l}^{b}$ is much tighter than $\Delta_{l}^{a}$ at the first step of the algorithm.
-Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice.
-The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
 5 Decentralized version
@@
 Computation $O(log^{2}(s)+n^{2}\cdot log(s))$ $O(s\cdot log^{2}(s)+sn^{2}\cdot log(s))$
-In this section, we explore the scenario where the matrix $\bm{A}$ is distributed across multiple clients.
-Specifically, each client holds a matrix $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$ such that $\sum_{i=1}^{s}\bm{A}^{(s)}=\bm{A}$.
-The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(s)}$ to the server or other clients.
-The randomized power method involves linear operations, which makes it well-suited for parallelization.
-An implementation of a private and distributed power method, using public channel communication between the clients and the server, was proposed in (Balcan et al., 2016).
-However, this approach relies on local Differential Privacy, since the data exchanged between the clients and the server can be observed by everyone.
-To improve privacy while keeping the benefits of distributed computation, we propose using Secure Aggregation, a lightweight Multi-Party Computation (MPC) protocol.
-Secure Aggregation eliminates the need for local DP by allowing us to use distributed DP, which offers privacy guarantees similar to central DP.
-This approach has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
-In distributed DP, each client adds noise to its own contributions before using the Secure Aggregation protocol.
-The noise is calibrated such that the sum of the noise across clients has a variance comparable to that used in Central DP, thus achieving similar privacy guarantees without requiring a trusted curator to aggregate the data.
-We assume the “honest-but-curious” threat model, meaning that while the clients may attempt to learn additional information, they follow the protocol honestly.
-We also assume no corrupted or dropout users.
-For simplicity, this paper focuses on a basic version of distributed DP, neglecting effects of data quantization or errors introduced by modular arithmetic.
-We refer to (Kairouz et al., 2021) for an implementation of Distributed DP that takes these effects into account.
-Finally, it is also possible to use a fully decentralized and differentially private protocol to perform averaging (Sabater et al., 2022).
-It is therefore straightforward to transform our protocol to avoid relying on a central server for Secure Aggregation, while keeping similar utility and communication costs.
+In this section, we consider a decentralized setting where the matrix $\bm{A}$ is distributed across multiple clients.
+Specifically, each client $i$ holds a private matrix $\bm{A}^{(i)}\in\mathbb{R}^{n\times n}$, such that the global matrix is the sum of these local matrices: $\bm{A}=\sum_{i=1}^{s}\bm{A}^{(i)}.$
+The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(i)}$ to the server or to other clients.
+The randomized power method involves linear operations, making it well-suited for parallelization and distributed computation.
+Balcan et al.
+(2016) proposed a private and federated power method using communication over public channels between clients and a server.
+However, this approach relies on Differential Privacy, since the data exchanged can be observed by everyone, which requires high levels of noise to ensure privacy.
+To enhance privacy while retaining the benefits of distributed computation, we use Secure Aggregation, a lightweight Multi-Party Computation protocol.
+Secure Aggregation allows clients to collaboratively compute sums without revealing individual data, enabling the use of distributed DP.
+This approach offers privacy guarantees similar to central DP and eliminates the need for a trusted curator.
+Distributed DP has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
+In distributed DP, each client adds carefully calibrated noise to their local contributions before participating in the Secure Aggregation protocol.
+The noise is designed such that the sum across clients has a variance comparable to that used in central DP, thereby achieving similar Differential Privacy guarantees without requiring a trusted aggregator, and allowing clients to keep the data locally.
+We operate under the “honest-but-curious“ threat model, where clients follow the protocol correctly but may attempt to learn additional information from received data.
+We also assume there are no corrupted or dropout users during the computation.
+For simplicity, this paper focuses on a basic implementation of distributed DP, neglecting the effects of data quantization or errors introduced by modular arithmetic.
+We refer to Kairouz et al.
+(2021) for an implementation of Distributed DP that takes these effects into account.
+Finally, it is possible to use a fully decentralized and differentially private protocol to perform averaging without relying on a central server for Secure Aggregation, as demonstrated by Sabater et al.
+(2022).
+Adopting such approaches can further enhance decentralization while maintaining similar utility and communication costs.
 5.1 Decentralized and private power method using distributed DP:
-We give a federated version of the PPM in Algorithm 2.
-In this version, we reduce the variance of the noise by a multiplicative factor in O($s\cdot p\log({d})\cdot\|\bm{U}\|_{\infty}^{2}$) compared to (Balcan et al., 2016) due to the fact that we emulate Central DP with Secure Aggregation (thus using a secure channel) rather than using a public channel that only allows for local DP, and because we propose a generally tighter sensitivity bound.
-We show the equivalence between Algorithm 2 and Algorithm 1 in Theorem 6.2, and provide the proof in Appendix A.5.
-We first define the secure aggregation algorithm.
+We introduce a federated version of the Privacy-Preserving Power Method in Algorithm 2.
+This version reduces the noise variance by a factor of $O(sp\log d\,\|\bm{U}\|_{\infty}^{2})$ compared to the method in Balcan et al.
+(2016).
+The improvement comes from the fact that we emulate centralized Differential Privacy (DP) over secure channels, avoiding the higher noise required in local DP settings with public channels, and that we use our generally tighter sensitivity bound.
+We demonstrate the equivalence between the decentralized Algorithm 2 and the centralized Algorithm 1 in Theorem 5.2 (proof provided in Appendix A.5).
+We also give a simplified definition of the Secure Aggregation algorithm.
 Definition 5.1.
@@
 Algorithm 2 Federated private power method
-Theorem 6.2 (Privacy and convergence of Algorithm 2).
-Algorithm 2 is as privacy-preserving and has the same utility as its centralized version (Algorithm 1).
+Theorem 5.2 (Privacy and utility of Algorithm 2).
+The decentralized Privacy-Preserving Power Method (Algorithm 2) provides the same privacy guarantees and achieves equivalent utility (in terms of convergence) as its centralized version (Algorithm 1).
 5.2 Communication and computation cost analysis:
@@
 Table 1 presents the communication and computations cost overhead introduced by the polylogarithmic version of Secure Aggregation that we employ (Bell et al., 2020).
-7 Conclusion
+6 Empirical comparison of the proposed bounds
+We introduced both runtime-dependent and runtime-independent convergence bounds for our algorithm.
+The runtime-dependent bound depends on the matrix $\bm{A}$ and the iteratively computed matrices $\bm{X}_{l}$, whereas the runtime-independent bound (Theorem 4.1) only depends on $\bm{A}$.
+To illustrate the practical impact of our changes, we focus on an application in recommender systems and compare how our algorithm perform in this context.
+We will then use a statistical approximation to see how the runtime-dependent bounds compare at the first step of the algorithm, regardless of the application.
+6.1 Application to recommender systems
+State-of-the-art recommender systems like GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) utilize singular value decomposition (SVD) or eigenvalue decomposition as part of their algorithms.
+Specifically, they represent the dataset (comprised of user-item interactions) using a normalized adjacency matrix $\tilde{\bm{R}}$, from which they compute a normalized item-item matrix $\tilde{\bm{P}}$.
+To get rid of noise in $\tilde{\bm{R}}$, they compute an ideal low-pass filter based on the top-$p$ eigenvectors of $\tilde{\bm{P}}$ and apply it to $\tilde{\bm{R}}$.
+Detailed definitions are provided in Appendix A.7.
+Table 2 presents key statistics of popular recommendation datasets used in the literature.
+Depending on the use case, these datasets may be held by a central curator, or distributed among the users of a recommender system.
+Accordingly, either the centralized PPM (Algorithm 1) or the decentralized PPM (Algorithm 2) can be applied to compute the leading eigenvectors.
+To demonstrate the practicality of our proposed methods, we focus on a decentralized setting where each user has access to part of the dataset and collaborate with other users to compute the desired eigenvectors.
+Table 2: Statistics of datasets
+Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{prior}$ (existing) ${\hat{\Delta}_{l}}$ (ours)
+Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.12 ×$\times\sqrt{p}$ 1
+MovieLens 71,567 10,677 7,972,582 0.39 32.84 1.19 ×$\times\sqrt{p}$ 1
+EachMovie 74,425 1,649 2,216,887 0.49 21.89 1.33 ×$\times\sqrt{p}$ 1
+Jester 54,906 151 1,450,010 0.73 7.78 1.64 ×$\times\sqrt{p}$ 1
+6.1.1 Empirical comparison of the runtime-independent bounds
+As discussed in Section 4.1, our proposed runtime-independent bound is tighter than those proposed in Hardt & Price (2014); Balcan et al.
+(2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$.
+Table 2 shows that this condition holds for popular recommender system datasets.
+As we can deduce from Table 2, the proposed bound theoretically allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
+6.1.2 Practical utility of the proposed algorithms
+We saw that the proposed runtime-independent bounds were practically tighter than the previous one for the task of interest.
+We demonstrate that our decentralized PPM can privately compute the top-$p$ eigenvectors $\bm{U}_{p}$ of the user-normalized item-item matrix $\tilde{\bm{P}}$ (proof in Appendix A.1):
+Lemma 6.1.
+Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ can compute $\bm{U}_{p}$, the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting under $(\epsilon,\delta)$-Differential Privacy.
+Figure 1: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the EachMovie dataset.
+Both GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) use $\bm{U}_{p}$ to compute the ideal low pass filter and apply it to the interaction matrix $\bm{R}$, giving a filtered interaction matrix $\bm{R}_{p}$.
+To illustrate the practicality of our proposed method, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using either our proposed decentralized PPM, or the PPM versions proposed by  Balcan et al.
+(2016).
+We then use $\tilde{\bm{U}}_{p}$ to compute an approximation of $\bm{R}_{p}$ (noted $\tilde{\bm{R}_{p}}$), and compute the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to each version.
+Figure 1 illustrates the impact of the privacy parameter $\epsilon$ on the approximation error for the EachMovie dataset, with $p=32$ and $L=3$.
+Our decentralized PPM achieves a relatively low approximation error (approximately 1/10 of the random guessing error) for practical values of $\epsilon$ (5-10).
+In contrast, the methods from Balcan et al.
+(2016) require $\epsilon$ values of at least $10^{3}$ to achieve comparable errors, which does not provide meaningful privacy protection.
+Additional comparisons and details are provided in Appendix A.7.1.
+6.2 Matrix-agnostic comparison of the runtime-dependent bounds
+We analyze the tightness of our proposed runtime-dependent bound compared to the previous bound in a matrix-agnostic manner, focusing on the first iteration of the algorithm.
+At iteration $l=0$, the leading singular vectors are initialized as $\bm{X}^{0}=\text{Q}(\bm{\Omega})$, where $\bm{\Omega}\sim\mathcal{N}(0,1)^{n\times k}$ and $\text{Q}(\bm{\Omega})$ denotes the orthonormal matrix obtained from the QR decomposition of $\bm{\Omega}$.
+Since $\bm{X}^{0}$ is independent of the matrix $\bm{A}$, it allows us to assess the relative tightness of the proposed bound $\hat{\Delta}_{0}$ compared to the previous bound $\Delta_{0}^{prior}$ at the first step.
+Let $r(k,n)$ be the ratio of the expected values of the two bounds, as a function of the desired number of eigenvectors $k$ and the matrix dimension $n$:
+$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{prior}]}{\mathbb{E}[\hat{\Delta}_{0}]}.$ (18)
+6.2.1 Theoretical approximation of $r(k,n)$:
+We derive an asymptotic approximation of $r(k,n)$ using Approximation 6.2 (detailed in Appendix A.4).
+Approximation 6.2.
+[…]
+$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
+$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
+6.2.2 Empirical approximation of $r(k,n)$:
+Figure 2: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=8000$ and $k$ ranging between 64 and 4000, with a step of 64.
+Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm.
+The theoretical ratios are estimated using Theorem 6.2.
+To compare the tightness of our proposed bound $\hat{\Delta}_{0}$ to the previous bound $\Delta_{0}^{prior}$, we perform both theoretical and empirical analyses at the first iteration, where the noise scaling depends only on the random initialization $\bm{X}^{0}$.
+Since $\bm{X}^{0}$ is independent of $\bm{A}$, we can estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$ by sampling random Gaussian matrices $\bm{X}^{0}$.
+Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$.
+We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$.
+We use Theorem 6.2 to approximate it theoretically.
+To measure it empirically, we compute $\Delta_{0}^{prior}$ and $\hat{\Delta}_{0}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm.
+We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$.
+We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples.
+We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 6.2.
+Figure 2 and Figure 3 (in Appendix) presents the comparison between empirical (blue dots) and theoretical (green line) estimates of $r(k,n)$ for $n=8000$ and $n=12000$.
+The empirical ratios are calculated from the averages over the $t=5$ runs for each value of $k$.
+The results indicate that both empirically and theoretically the proposed noise scaling $\hat{\Delta}_{l}$ is much tighter than $\Delta_{l}^{prior}$ at the first step of the algorithm.
+Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice.
+The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
+7 Conclusion and limitations
 We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems.
 We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees.
-By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, leading to higher utility without compromising on privacy.
-The decentralized variant of the algorithm maintains the efficiency and privacy of the centralized version but adapts it for distributed environments.
-Our theoretical and empirical results demonstrate the utility and privacy advantages of our approach, suggesting that it offers a strong solution for real-world decentralized applications.
-8 Limitations
-The proposed method is effective in ensuring differential privacy and decentralization but still has limitations.
-The privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic in certain environments.
+By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, maintaining the efficiency and privacy of the centralized version but adapting it for distributed environments.
+However, our privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic.
 In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
+8 Impact Statement (required by ICML)
+Our contributions directly address the need for privacy-preserving algorithms in the context of large-scale machine learning systems, where sensitive personal or organizational data is often at stake.
+The decentralized implementation ensures that the method can be applied in some real-world environments where data is distributed across multiple users or devices, making it particularly suitable for recommender systems and other distributed applications.
+This work is expected to have implications for privacy-conscious machine learning in real-world systems, to perform large-scale data analysis while maintaining robust privacy protections.
 Appendix A Proofs of Results
@@
 By injecting (22) into (23), we have:
-$\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\Delta_{l}^{b}.\\ \end{split}$ (24)
-Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\Delta_{l}^{b}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
-A.2 Proof of Theorem 3.2
+$\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\hat{\Delta}_{l}.\\ \end{split}$ (24)
+Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\hat{\Delta}_{l}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
+A.2 Privacy proof
+A.2.1 Note on related privacy proofs:
+Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by Hardt & Price (2014), Balcan et al.
+(2016), Hardt & Roth (2013), Hardt & Roth (2012).
+These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP.
+Specifically, Balcan et al.
+(2016) references the privacy proof from Hardt & Price (2014), which builds upon the privacy proof from Hardt & Roth (2013).
+The proof therein is also closely related to the one of Hardt & Roth (2012).
+However, both Hardt & Roth (2013) (Theorem 2.4) and Hardt & Roth (2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped.
+This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all.
+The original composition rule is presented in Theorem III.3 of Dwork et al.
+(2010).
+Moreover, Lemma 3.4 of Hardt & Roth (2013) misuses Theorem 2.4.
+Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration.
+By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
+A.2.2 Proof of Theorem 3.2
 Proof.
@@
 $\rho$ $=\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$ (25)
-$\leq\sum_{i=1}^{L}\frac{(\Delta_{l}^{b})^{2}}{2\sigma_{i}^{2}}$ (26)
+$\leq\sum_{i=1}^{L}\frac{(\hat{\Delta}_{l})^{2}}{2\sigma_{i}^{2}}$ (26)
 $=\sum_{i=1}^{L}\frac{1}{2\sigma^{2}}$ (27)
@@
 $=\frac{\epsilon^{2}}{{8\log(1/\delta)}}.$ (29)
-By Proposition 1.3 of (Bun & Steinke, 2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$.
+By Proposition 1.3 of Bun & Steinke (2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$.
 Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
@@
 $\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}\,.$ (41)
-Which leads us to the following Theorem by injecting $\sigma_{l}=\Delta_{l}^{b}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
-A.4 Derivation of Approximation 5.1
-We provide here the derivation for Approximation 5.1:
+Which leads us to the following Theorem by injecting $\sigma_{l}=\hat{\Delta}_{l}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
+A.4 Derivation of Approximation 6.2
+We provide here the derivation for Approximation 6.2:
 Proof.
@@
 By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$.
-Then by Theorem 2.2.1 of (Chikuse, 2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries.
+Then by Theorem 2.2.1 of Chikuse (2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries.
 We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty.
@@
 $k\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k^{2}\sigma^{2}}{n^{2}})$ $\sum_{j}\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k\sigma^{2}}{n^{2}}).$ (45)
-We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 of (Massart, 1896).
+We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 from Massart (1896).
 We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$.
-We assume that the variances of $\Delta_{l}^{a}$ and $\Delta_{l}^{b}$ (as $n$ and $k$ are large) are small and therefore:
-$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\mathbb{E}[(\Delta_{l}^{a})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
-$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\mathbb{E}[(\Delta_{l}^{b})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
+We assume that the variances of $\Delta_{l}^{prior}$ and $\hat{\Delta}_{l}$ (as $n$ and $k$ are large) are small and therefore:
+$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\mathbb{E}[(\Delta_{l}^{prior})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
+$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\mathbb{E}[(\hat{\Delta}_{l})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
+Figure 3: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=12000$ and $k$ ranging between 64 and 4000, with a step of 64.
+Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm.
+The theoretical ratios are estimated using Theorem 6.2.
 A.5 Proof of Theorem 4.1
@@
 Proof.
-First bound: We know by Section 2.4 of (Woodruff et al., 2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1.
+First bound: We know by Section 2.4 from Woodruff et al.
+(2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1.
 Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
 Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM.
-As proposed in (Hardt & Price, 2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar.
+As proposed in Hardt & Price (2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar.
 Therefore we have:
@@
 where $\bm{G}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\cdot(s\nu^{2}))^{n\times p}$, and we have $s\nu^{2}=\sigma^{2}$ by definition, completing the equivalence proof. ∎
+A.7 Application to recommender systems:
+Let $s$ be the number of users in our system and $n$ the number of items.
+Let $\bm{R}\in\mathbb{R}^{s\times n}$ be the user-item interaction matrix, such that $\bm{R}_{ui}=1$ only if user $u$ has interacted with item $i$, and $\bm{R}_{ui}=0$ else.
+Let $\bm{U}=\text{Diag}(\bm{R}\cdot\bm{1}_{|\mathcal{I}|})$ be the user degrees matrix, and $\bm{I}=\text{Diag}(\bm{1}_{|\mathcal{I}|}^{\top}\cdot\bm{R})$ the item degrees matrix.
+In Shen et al.
+(2021); Choi et al.
+(2023), the normalized interaction matrix is defined as:
+$\bm{R}^{\prime}=\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}},$
+and the item-item normalized adjacency matrix as:
+$\begin{split}\bm{P}^{\prime}&=\tilde{\bm{R}}^{\top}\tilde{\bm{R}}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}}).\\ \end{split}$ (58)
+To simplify our analysis, we consider $\bm{I}$ public and do not use item-wise normalization in the computation of ideal low pass filter, leaving it for future work.
+We therefore define $\tilde{\bm{R}}=\bm{U}^{-\frac{1}{2}}\bm{R}$ and $\tilde{\bm{P}}=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})$.
+Lemma A.1.
+We can use Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ to compute the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting with under a $(\epsilon,\delta)$-Differential Privacy guarantee.
+Proof.
+We note $\tilde{\bm{P}}_{ij}$ as the element of $\tilde{\bm{P}}$ at line $i$ and column $j$ and $d_{user}(u)=\sum_{i=0}^{n-1}r_{ui}$. We can write $\tilde{\bm{P}}_{ij}$ as such:
+$\begin{split}\tilde{\bm{P}}_{ij}&=((\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}))_{ij}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{T})_{i,*}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=((\bm{U}^{-\frac{1}{2}}\bm{R})_{*,i})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=\sum_{u=0}^{s-1}\frac{1}{\sqrt{d_{user}(u)}}r_{ui}\frac{1}{\sqrt{d_{user}(u)}}r_{uj}\\ &=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}.\\ \end{split}$ (59)
+By noticing that $(\bm{R}^{\top}\bm{R})_{ij}=\sum_{u}r_{ui}\cdot r_{uj}$, then we can deduce that $\tilde{\bm{P}}=\sum_{u}\frac{1}{d_{user}(u)}\bm{R}_{u}^{\top}\bm{R}_{u}$.
+Therefore $\tilde{\bm{P}}$ is partitioned among $s$ users as described in Section 5.
+Sensitivity: We protect the user at the item-level and use the deletion model of differential privacy to compute the sensitivity of the PPM used with the item-item normalized adjacency matrix ($\bm{A}=\tilde{\bm{P}}$).
+Therefore we have:
+$\bm{A}_{ij}=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj},$ (60)
+and
+$\bm{A}^{\prime}_{ij}=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)-1}\cdot r_{ui}^{\prime}\cdot r_{uj}^{\prime},$ (61)
+where $r_{ui}^{\prime}=r_{ui}$ except for one user-item interaction, i.e.
+$r_{vk}=1$ but $r_{vk}^{\prime}=0$.
+Let $\bm{C}=\bm{A}-\bm{A}^{\prime}$.
+Let $\mathcal{N}(v)$ be the set of items which user $v$ liked before deletion.
+We have:
+$\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}$ $=\sum_{i}(\sum_{j}|\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}-\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)-1}\cdot r_{ui}^{\prime}\cdot r_{uj}^{\prime}|)^{2}$ (62)
+$=\sum_{i}(\sum_{j}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (63)
+$=\sum_{i\in\mathcal{N}(v)}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (64)
+$=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (65)
+$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}.$ (66)
+We have:
+$\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (68)
+$=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\{\mathcal{N}(v)\backslash k\}}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|+\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vk})^{2}$ (69)
+$=(d_{user}(v)-1)\{(d_{user}(v)-1)|\frac{1}{d_{user}(v)}-\frac{1}{d_{user}(v)-1}|+\frac{1}{d_{user}(v)}\}^{2}$ (70)
+$=(d_{user}(v)-1)\{|\frac{1}{d_{user}(v)}|+\frac{1}{d_{user}(v)}\}^{2}$ (71)
+$=\frac{2(d_{user}(v)-1)}{d_{user}(v)^{2}},$ (72)
+and
+$(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (74)
+$=(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vj}|)^{2}$ (75)
+$=(d_{user}(v)|\frac{1}{d_{user}(v)}|)^{2}$ (76)
+$=1.$ (77)
+By noticing that $\frac{2(d_{user}(v)-1)}{d_{user}(v)^{2}}\leq 1$, we can deduce that:
+$\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}$ $=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (79)
+$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (80)
+$=\frac{2(d_{user}(v)-1)}{d_{user}(v)^{2}}+1$ (81)
+$\leq 2$ (82)
+$\implies$ $\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\leq\sqrt{2}.$ (83)
+By Equation 24,
+$\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}$ (84)
+$\implies\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{2}.$ (85)
+A.7.1 Approximation errors comparisons:
+As explained in Section 6.1.2, GF-CF and BSPM (with its parameter $T_{b}=1$) use $\bm{U}_{p}$ to compute the ideal low pass filter and filter the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$.
+Indeed, we have:
+$\bm{R}_{p}=\bm{R}\cdot\bm{I}^{-\frac{1}{2}}\bm{U}_{p}\bm{U}_{p}^{\top}\bm{I}^{\frac{1}{2}}.$ (86)
+To study the impact of Differential Privacy on our system, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using our proposed decentralized PPM, or the PPM versions of Balcan et al.
+(2016).We have:
+$\tilde{\bm{R}}_{p}=\bm{R}\cdot\bm{I}^{-\frac{1}{2}}\tilde{\bm{U}}_{p}\tilde{\bm{U}}_{p}^{\top}\bm{I}^{\frac{1}{2}}.$ (87)
+We then define the relative approximation error caused by the use of differential privacy as $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$.
+We use $L=3$ to run the Power Method as it is the default hyper-parameter choice from Shen et al.
+(2021); Choi et al.
+(2023).
+We use $p=32$ to have acceptable approximation error (on the order of 0.02-0.05 for numbers bounded between 0 and 1) for reasonable values of $\epsilon$ (5-10).
+We showed the impact of the Differential Privacy parameter $\epsilon$ on the approximation errors for the EachMovie dataset in Figure 1.
+We get similar results for the MovieLens and Jester datasets in Figures 5 and  4.
+Figure 4: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the Jester dataset.
+Figure 5: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the MovieLens dataset.

### v3

@@
 The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks.
-However, its application to large datasets containing personal information (e.g.
-web interactions, search history, personal tastes) raises critical privacy concerns.
-This paper addresses these issues by proposing enhanced privacy-preserving variants of the method, which reduce the variance of the noise introduced to achieve Differential Privacy (DP).
-We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy.
-We leverage Secure Aggregation to run the algorithm on data distributed among multiple users or devices, while using a noise scale in the decentralized setting that is similar to the one in the centralized setting yet without revealing individual data.
-We improve upon existing convergence bounds for both the centralized and decentralized versions.
-Machine Learning, ICML
+However, its application to large datasets containing personal information (e.g., web interactions, search history, personal tastes) raises critical privacy problems.
+This paper addresses these issues by proposing enhanced privacy-preserving variants of the method.
+First, we propose a variant that reduces the amount of the noise required in current techniques to achieve Differential Privacy (DP).
+More precisely, we refine the privacy analysis so that the Gaussian noise variance no longer grows linearly with the target rank, achieving the same $(\varepsilon,\delta)$-DP guarantees with strictly less noise.
+Second, we adapt our method to a decentralized framework in which data is distributed among multiple users.
+The decentralized protocol strengthens privacy guarantees with no accuracy penalty and a low computational and communication overhead.
+Our results include the provision of tighter convergence bounds for both the centralized and decentralized versions, and an empirical comparison with previous work using real recommendation datasets.
 1 Introduction
-The randomized power method is a simple and computationally efficient tool for solving large-scale linear algebra problems, particularly in spectral analysis and recommendation systems.
-It approximates the principal singular vectors of a matrix with low computational complexity (nearly linear), making it particularly suitable for analyzing large modern datasets.
-In large-scale machine learning systems, protecting user privacy is paramount.
-Datasets often contain sensitive personal or organizational information, so privacy has become a critical requirement.
-The standard randomized power method does not inherently provide privacy guarantees.
-While its output might seem less sensitive than the input data, there is no formal guarantee that private information embedded in the data cannot be inferred.
-To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides strong formal guarantees and mitigate potential privacy leaks of an algorithm.
-DP ensures that the output of an algorithm reveals little about any individual record in the input.
+The randomized power method has emerged as an efficient and scalable tool for addressing large-scale linear algebra problems central to modern machine learning pipelines.
+By constructing an orthonormal basis for a matrix’s range in near-linear time, the method scales seamlessly to practical large datasets.
+Its reliance on simple matrix products ensures compatibility with sparse data representations and enables efficient parallelization and hardware acceleration on GPUs and distributed architectures.
+Beyond its simplicity, the method provides strong approximation guarantees and accelerates a wide spectrum of applications.
+It has been used for principal component analysis (PCA) (Journée et al., 2010), singular value decomposition (SVD) (Halko et al., 2011), truncated eigen-decompositions (Yuan and Zhang, 2013), and matrix completion (Feng et al., 2018).
+Extensions have powered recommender systems (e.g., Twitter (Gupta et al., 2013), GF-CF (Shen et al., 2021), BSPM (Shen et al., 2021)), PageRank-style ranking (Ipsen and Wills, 2005), PDE solvers (Greengard and Rokhlin, 1997), or large-scale least-squares and linear-system solvers (Rokhlin and Tygert, 2008).
+However, as these methods are integrated into large-scale machine learning systems involving personal data, protecting user privacy is paramount.
+Unfortunately, the standard randomized power method does not inherently provide privacy guarantees.
+While its output might seem less sensitive than the input data, there is no formal guarantee against inference of private information embedded in the data.
+To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful framework that provides strong guarantees and mitigates potential privacy leaks of an algorithm, ensuring that the output of an algorithm reveals little about any individual record in the input.
 Several works have attempted to apply DP to the randomized power method.
-For instance, Hardt & Price (2014) and Balcan et al.
-(2016) introduced differentially private versions of the method, and Guo et al.
-(2021); Wang & Xu (2020) explored federated approaches to private power iteration.
-Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
+For example, Hardt and Price (2014); Balcan et al.
+(2016) developed centralized DP variants of the power method, whereas Wang and Xu (2020); Guo et al.
+(2024) investigated federated DP protocols that can be used when data is kept locally across multiple devices.
+Adjacent works explore DP versions of PCA in both centralized (Liu et al., 2022) and federated (Wang and Xu, 2020; Briguglio et al., 2023) settings and give optimal convergence bounds under distributional assumptions.
 Despite these advancements, existing approaches suffer from several limitations.
-First, their performance heavily depends on the number of singular vectors being computed (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021; Liu et al., 2022; Wang & Xu, 2020), which impacts both utility and privacy guarantees.
-Second, they are primarily designed for centralized settings (Hardt & Price, 2014; Liu et al., 2022), where a trusted curator is assumed to hold the data.
-Moreover, some methods (Liu et al., 2022) make strong assumptions about the underlying data distribution (i.e.
-that the data is sub-Gaussian) which makes it harder to use these methods in practice.
-Some federated versions (Briguglio et al., 2023) claim to guarantee privacy by use of the federated setting, although it has been shown that decentralization does not offer privacy by design (Geiping et al., 2020).
-Finally, no fully decentralized versions exist to our knowledge.
-The previously proposed federated versions (Balcan et al., 2016; Guo et al., 2021; Wang & Xu, 2020) use public channel communication and ignore advances in Multi-Party-Computation (MPC).
-This makes them unsuitable for decentralized environments, such as distributed recommender systems and social networks, where data is partitioned across users and/or devices and communications are restricted to a predefined communication graph.
+First, their performance heavily depends on the number of singular vectors being computed (Hardt and Price, 2014; Balcan et al., 2016; Guo et al., 2024; Liu et al., 2022; Wang and Xu, 2020), which impacts both utility and privacy guarantees.
+Second, they are designed for centralized settings (Hardt and Price, 2014; Liu et al., 2022), where a trusted curator is assumed to hold the data.
+Moreover, some methods (Grammenos et al., 2020; Liu et al., 2022) make strong assumptions about the data distribution (e.g., sub-Gaussianity) which makes it harder to use these methods in practice.
+Some federated versions (Briguglio et al., 2023; Hartebrodt et al., 2024) claim to guarantee privacy due to the federated setting, but it has been shown that decentralization does not offer privacy by design (Geiping et al., 2020).
+MOD-SuLQ (Chaudhuri et al., 2013) and its federated and streaming PCA variants (Grammenos et al., 2020) offer strong guarantees but are specifically tailored for reconstructing the top principal component ($k=1$).
+These methods add noise directly to the covariance matrix, rely on direct, computationally costly exact singular value decompositions (SVD) and are tailored for settings for which the number of samples largely exceeds the dimensionality ($n\gg d$).
+In contrast, the randomized power method iteratively adds noise directly to the approximation of the singular vectors themselves rather than the input matrix.
+This strategy reduces the dependence on the dimensionality of the original data, making it computationally more efficient and scalable.
+Memory-limited, streaming PCA methods such as those proposed by Mitliagkas et al.
+(2013) are optimized for sequential processing under memory constraints but lack privacy guarantees, requiring additional modifications.
+Finally, no fully decentralized versions exist to our knowledge, making them unsuitable for decentralized environments (e.g., recommender systems and social networks), where data is partitioned across users/devices and communications are restricted to a predefined communication graph.
+The previously introduced DP federated versions (Balcan et al., 2016; Wang and Xu, 2020; Guo et al., 2024) use public channel communication and therefore require local DP, which hinders convergence.
 1.1 Contributions
-The core focus of this article is to introduce improved privacy-preserving randomized power methods and to extend them to a decentralized setting, while maintaining computational efficiency and performance.
-We propose new convergence bounds for the (decentralized) privacy-preserving power method; these show that our proposed algorithm has a weaker dependence on the number of singular vectors that are computed.
-We propose an improved differentially private version of the randomized power method, with tighter convergence bounds, reducing the noise required to achieve privacy and therefore improving the accuracy of the method.
-In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method.
-[^1] This proof addresses some minor mistakes in a privacy proof from Hardt & Roth (2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016).
-The modified result allows us to use a wider range of privacy parameters.
-Finally, we propose decentralized versions of the algorithms.
-These are particularly relevant in contexts where data is distributed across multiple users or devices, as is often the case in recommender systems.
-To address the privacy concerns of the users, without forcing them to rely on a trusted curator, we propose a federated privacy-preserving version of the method using Secure Aggregation (Bell et al., 2020), a form of Multi-Party Computation (MPC) and a fully decentralized version using GOPA (GOssip noise for Private Averaging) (Sabater et al., 2022).
-This allows users to keep their data local while contributing to a privacy-preserving global computation.
-Our decentralized approach preserves the effectiveness of the centralized version of the randomized power method while incorporating the privacy advantages of local differential privacy (DP).
-Notably, by employing Secure Aggregation, we are able to set the noise scale comparable to that of central DP, enhancing accuracy.
-This makes the proposed method a promising candidate for decentralized and privacy-conscious applications.
-2 Background Material
-2.1 Notation:
-Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a real-valued positive semi-definite matrix, and let $k$ denote the target rank, where typically $k\ll n$.
-We also define $\eta$ as a small positive scalar that represents the precision of the approximation.
-The eigenvalue decomposition of $\bm{A}$ is given by:
-$\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$ (1)
-where $\bm{U}$ is the matrix of eigenvectors of $\bm{A}$, and $\bm{\Lambda}$ is a diagonal matrix containing the corresponding eigenvalues on its diagonal.
-Let $\mathcal{N}(\mu,\sigma^{2})^{n\times p}$ be a ($n\times p$) random matrix of which each element is a random variable independent and identically distributed (i.i.d.) according to a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
-For any matrix $\bm{X}$, let its element-wise maximum norm be $\|\bm{X}\|_{\max}=\max_{i,j}\|\bm{X}_{ij}\|$, where $\bm{X}_{ij}$ is the $(i,j)$-th entry of matrix $\bm{X}$.
-We denote by $\bm{X}_{j:}$ the $j$-th row of $\bm{X}$.
-We also define two coherence measures for the matrix $\bm{A}$:
+In this work, we propose a low-noise, privacy-preserving variant of the randomized power method whose approximation error remains bounded regardless of subspace dimension, significantly improving its applicability to practical datasets.
+We additionally propose a decentralized variant that strengthens privacy guarantees while maintaining the computational efficiency, and approximation guarantees of the centralized method.
+We summarize our contributions below:
+  • An Improved Differentially Private Randomized Power Method.
+We introduce a Differentially Private variant of the randomized power method under a generalized adjacency model, making it more suitable for practical applications.
+Unlike prior works, which relied on loose sensitivity bounds extrapolated from one-dimensional vectors, we derive a significantly tighter sensitivity analysis directly applicable to the multidimensional case.
+We show theoretically with refined convergence bounds and empirically with a recommender system use-case [^1] We provide an anonymous code repository here.
+that this refinement considerably reduces the amount of noise and the approximation error for a given privacy budget.
+  • Decentralized private power method.
+We propose a decentralized version of our algorithm, allowing settings for which data is distributed across multiple users or devices (e.g., recommender systems).
+This version is compatible with Secure Aggregation (Bell et al., 2020; Bonawitz et al., 2017; Kadhe et al., 2020), or fully decentralized private averaging approaches using correlated noise (Sabater et al., 2022; Allouah et al., 2024).
+Our decentralized approach preserves the effectiveness of the centralized version while incorporating Distributed Differential Privacy.
+  • Rigorous privacy guarantees.
+We present a new, more straightforward proof of a privacy result for the differentially private randomized power method.
+This proof addresses some mistakes in a privacy proof from Hardt and Roth (2012, 2013) that have been reproduced in several follow-up works (Hardt and Price, 2014; Balcan et al., 2016) and allows for a wider range of DP parameters.
+The remainder of the paper is organized as follows: In Section 2 we review the necessary background and notation.
+Section 3 introduces our generalized adjacency notion, derives the improved sensitivity bound (Theorem 3.1), and presents the overall privacy proof (Theorem 3.2) together with runtime-dependent convergence guarantees (Theorem 3.3).
+In Section 4 we turn this into a fully runtime-independent bound (Theorem 4.1).
+Section 5 develops the decentralized variant (Algorithm 2), proves its equivalence to the centralized version (Theorem 5.2), and analyzes its communication and computation overhead.
+We empirically compare the methods on standard recommendation datasets in Section 6.1, and conclude in Section 7 with a discussion of limitations and future directions.
+2 Background Material and Related Work
+Matrix Norms and Notations.
+For any matrix $\bm{X}$, the element-wise maximum norm is defined as $\|\bm{X}\|_{\max}=\max_{i,j}|\bm{X}_{ij}|$, where $\bm{X}_{ij}$ is the $(i,j)$-th element of $\bm{X}$.
+The $\ell_{2}$-norm is denoted as $\|\bm{X}\|_{2}$ and the Frobenius norm as $\|\bm{X}\|_{F}$.
+We use $\bm{X}_{j:}$ to denote the the $j$-th row of $\bm{X}$.
+Eigenvalue Decomposition.
+Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a real-valued positive semi-definite matrix, where $n$ is a positive integer.
+The eigenvalue decomposition of $\bm{A}$ is given by $\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$ where $\bm{U}\in\mathbb{R}^{n\times n}$ is a matrix of eigenvectors and $\bm{\Lambda}\in\mathbb{R}^{n\times n}$ is a diagonal matrix containing corresponding eigenvalues.
+QR Decomposition.
+We will use the matrix QR decomposition, obtained using the Gram-Schmidt procedure.
+Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, the QR decomposition factorizes it as $\bm{X}=\bm{Q}\bm{R}$, where $\bm{Q}\in\mathbb{R}^{n\times p}$ is an orthonormal matrix (i.e., $\bm{Q}^{\top}\bm{Q}=\bm{I}$) and $\bm{R}\in\mathbb{R}^{p\times p}$ is an upper triangular matrix.
+Gaussian Random Matrices.
+We denote by $\mathcal{N}(\mu,\sigma^{2})^{n\times p}$ a $(n\times p)$ random matrix where each element is an independent and identically distributed (i.i.d.) random variable according to a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
+Coherence Measures of a Matrix.
+We define below two coherence measures for the matrix $\bm{A}$, which will allow us to give convergence bounds for our method.
   • The $\mu_{0}$-coherence of $\bm{A}$ is the maximum absolute value of its eigenvector entries, defined as $\mu_{0}(\bm{A})=\|\bm{U}\|_{\max}=\max_{i,j}|\bm{U}_{ij}|$.
-  • The $\mu_{1}$-coherence of $\bm{A}$ is the maximum row sum of the absolute values of its eigenvector entries, defined as $\mu_{1}(\bm{A})=\|\bm{U}\|_{\infty}=\max_{i}(\sum_{j}|\bm{U}_{ij}|)$.
-Finally, we will also use the QR decomposition of matrix obtained using the Gram-Schmidt procedure.
-Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, the QR decomposition factorizes it as $\bm{X}=\bm{Q}\bm{R}$, where $\bm{Q}$ is an orthonormal matrix (i.e., $\bm{Q}^{\top}\bm{Q}=\bm{I}$) and $\bm{R}$ is an upper triangular matrix.
-2.2 Differential privacy:
-2.2.1 Approximate differential privacy:
-A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (adjacent datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
+  • The $\mu_{1}$-coherence of $\bm{A}$ is the maximum row $\ell_{2}$-norm of its eigenvectors, defined as $\mu_{1}(\bm{A})=\|\bm{U}\|_{2,\infty}=\max_{i}(\|\bm{U}_{i:}\|_{2})$.
+2.1 Differential privacy
+With positive integers $n$ and $m$ specifying the matrix dimensions, let $D_{1}\in\mathbb{R}^{n\times m}$ and $D_{2}\in\mathbb{R}^{n\times m}$ be two matrices representing two datasets embedding sensitive information.
+$D_{1}$ and $D_{2}$ are said to be adjacent ($D_{1}\sim D_{2}$) if they differ on one sensitive element of the dataset.
+For example, in a recommender system application, $D_{1}$ and $D_{2}$ can be binary user-item interaction matrices and a sensitive element of the dataset can be an entry in the matrix, corresponding to a user-item interaction.
+A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-Differential Privacy (DP) if for all adjacent datasets $D_{1}$ and $D_{2}$, and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
 $\Pr[\mathcal{M}(D_{1})\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(D_{2})\in S]+\delta,$ (1)
-where $\epsilon$ is a small positive parameter representing the privacy loss (smaller values indicate stronger privacy guarantees), and $\delta$ is a small probability representing the chance that the privacy guarantee may fail, allowing for a small probability of privacy breach.
-2.2.2 $\ell_{2}$-sensitivity:
-The $\ell_{2}$-sensitivity of a function $f$, denoted as $\Delta_{2}(f)$ or $\Delta_{2}$ for brevity, measures the maximum change in the output of $f$ when its input changes by at most one element (for neighboring datasets $D_{1}\sim D_{2}$).
-It is defined as:
-$\Delta_{2}=\max_{D_{1}\sim D_{2}}\|f(D_{1})-f(D_{2})\|_{2},$ (3)
-where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm.
-2.2.3 Zero-Concentrated Differential Privacy (zCDP):
-A randomized algorithm $\mathcal{M}$ is said to satisfy $\rho$-zero-Concentrated Differential Privacy (zCDP) if for all neighboring datasets $D_{1}$ and $D_{2}$, and for all $\alpha\in(1,\infty)$, the following holds:
-$D_{\alpha}(\mathcal{M}(D_{1})\|\mathcal{M}(D_{2}))\leq\rho\alpha\,,$ (4)
-where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees).
-The following lemma, introduced by Bun & Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
-Lemma 2.1 (Gaussian Mechanism (Proposition 1.6 (Bun & Steinke, 2016))).
-Let $f:X^{n}\to\mathbb{R}$ be a sensitivity-$\Delta$ function.
-Consider the mechanism (randomized algorithm) $M:X^{n}\to\mathbb{R}$, defined as $M(x)=f(x)+Z_{x}$ where for each input $x$, $Z_{x}$ is independently drawn from $\mathcal{N}(0,\sigma^{2})$.
-Then $M$ satisfies $\left(\frac{\Delta^{2}}{2\sigma^{2}}\right)$-zCDP.
-The next lemma, which is a generalized version of a result presented by Bun & Steinke (2016), explains how a randomized algorithm, constructed by recursively composing a sequence of zCDP-satisying randomized algorithms, also satisfies zCDP.
-Lemma 2.2 (Adaptive composition (Generalization from Lemma 2.3 (Bun & Steinke, 2016))).
-Let $M_{1}:X^{n}\to Y_{1}$, $M_{2}:X^{n}\times Y_{1}\to Y_{2}$, …, $M_{L}:X^{n}\times Y_{1}\times\dots\times Y_{L-1}\to Y_{L}$ be randomized algorithms.
-Suppose $M_{i}$ satisfies $\rho_{i}$-zCDP as a function of its first argument for each $i=1,2,\dots,L$.
-Let $M^{\prime\prime}:X^{n}\to Y_{L}$, constructed recursively by:
-$M^{\prime\prime}(x)=M_{L}(x,M_{L-1}(x,\dots,M_{2}(x,M_{1}(x))\dots)).$ (5)
-Then $M^{\prime\prime}$ satisfies $(\sum_{i=1}^{L}\rho_{i})$-zCDP.
-2.3 Privacy preserving power iteration:
-Let $k$ denote the target rank, $q$ a small real and $\eta$ a small positive value controlling the precision of the approximation.
-The goal of the randomized power method is to compute a matrix $\bm{Q}\in\mathbb{R}^{n\times p}$, where $p=k+q$, such that the range of $\bm{Q}$ approximates the range of the top $k$ eigenvectors of $\bm{A}$, which are denoted $\bm{U}_{k}$.
-Specifically, we aim to ensure that:
-$\|\bm{U}_{k}-\bm{Q}\bm{Q}^{\top}\bm{U}_{k}\|\leq\eta.$
-The privacy-preserving randomized power method extends this approach to protect sensitive information in $\bm{A}$ during computation.
-Prior work (Hardt & Price, 2014; Balcan et al., 2016) show that under approximate Differential Privacy (DP) parameters $(\epsilon,\delta)$, $\bm{Q}$ can be computed using Algorithm 1, where the sensitivity $\Delta_{l}$ at iteration $l$ is bounded by $\Delta_{l}^{\text{prior}}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$.
-Here, $\Delta_{l}^{\text{prior}}$ represents an $\ell_{2}$-sensitivity bound that quantifies the impact of a single-element change in $\bm{A}$ on $\bm{A}\bm{X}^{l-1}$.
-Specifically, $\Delta_{l}\leq\Delta_{l}^{\text{prior}}$.
-With this bound [^2] As defined in Fig.
-3 of (Hardt & Price, 2014), Alg.
-2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021)., the standard deviation of the noise $\sigma_{l}$ associated to DP at iteration $l$ is given by:
-$\sigma_{l}$ $=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$ (6)
-$=\Delta_{l}^{\text{prior}}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}.$ (7)
-As a first contribution, we refine this sensitivity analysis by introducing a significantly tighter bound for $\Delta_{l}$, reducing the noise required for DP and therefore allowing tighter convergence bounds (smaller $\eta$).
-1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$.
+where $\epsilon$ is a small positive scalar representing the privacy loss (smaller values indicate stronger privacy guarantees), and $\delta$ is a (typically negligible) probability that the privacy guarantee fails.
+Let $f:\mathbb{R}^{n\times m}\to\mathbb{R}^{d}$ be a query function associated with a mechanism $\mathcal{M}$.
+DP guarantees of mechanisms are defined using the sensitivity of $f$.
+In our contribution we use the $\ell_{2}$-sensitivity, denoted by $\Delta_{2}(f)$ or $\Delta_{2}$ and defined as $\Delta_{2}=\max_{D_{1}\sim D_{2}}\|f(D_{1})-f(D_{2})\|_{2}\,.$
+2.2 Privacy-Preserving Randomized Power Method
+Let $k$ be the target rank, let $b>0$ be a small positive integer, and let $\eta>0$ be an approximation tolerance.
+Given a positive semi-definite (PSD) matrix $\bm{A}\in\mathbb{R}^{n\times n}$, the aim of the randomized power method is to construct a matrix $\bm{Q}\in\mathbb{R}^{n\times p}$, where $p=k+b$, whose column space approximates that of the top $k$ eigenvectors of $\bm{A}$, i.e., $\bm{U}_{k}\in\mathbb{R}^{n\times k}$.
+Specifically, it aims to satisfy
+$\bigl{\|}\bm{U}_{k}-\bm{Q}\bm{Q}^{\top}\bm{U}_{k}\bigr{\|}\leq\eta.$ (2)
+To protect sensitive information in $\bm{A}$ while computing $\bm{Q}$, prior work (Hardt and Price, 2014; Balcan et al., 2016) shows that the randomized power method can be implemented as Algorithm 1 for it to satisfy $(\epsilon,\delta)$-Differential Privacy with adjacency defined as a single element change in $\bm{A}$.
+The corresponding adjacency notion is defined in more detail in Equation (6).
+To compute the standard deviation of the noise required to satisfy $(\epsilon,\delta)$-DP, one needs to bound the $\ell_{2}$-sensitivity $\Delta_{l}$ of the computations at each iteration of the algorithm.
+Prior works (Hardt and Price, 2014; Balcan et al., 2016) use the estimate
+$\Delta_{l}^{\text{prior}}\triangleq\sqrt{p}\|\bm{X}^{l}\|_{\max},$ (3)
+which upper-bounds the true sensitivity, i.e., $\Delta_{l}\leq\Delta_{l}^{\text{prior}}$. Here, $\Delta_{l}$ bounds the change on $\bm{A}\bm{X}^{l-1}$ under a single-element perturbation in $\bm{A}$.
+Using this bound [^2] See Fig.
+3 of Hardt and Price (2014), Alg.
+2 of Balcan et al.
+(2016), and Theorem 6 of Guo et al.
+(2024)., the standard deviation $\sigma_{l}$ of the Gaussian noise added at iteration $l$ to achieve $(\epsilon,\delta)$-DP is $\sigma_{l}=\sqrt{p}\,\bigl{\|}\bm{X}^{l}\bigr{\|}_{\max}\epsilon^{-1}\sqrt{4L\ln(1/\delta)}\,,$ where $L$ is the total number of power iterations.
+1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$
 2:  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
@@
 2.3 Existing convergence bounds
-To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in Balcan et al.
-(2016).
-In Theorem 2.4, we present a reduction of their Theorem 3.1.
-The theorem, and several others we present relies on the verification of the conditions in Assumption 2.3.
+To our knowledge, the strongest convergence bound for the privacy-preserving randomized power method (Algorithm 1) uses the bound $\Delta_{l}\leq\Delta_{l}^{prior}$ and is given by Balcan et al.
+(2016) in their Corollary 3.1.
+The following theorems rely on the conditions in Assumption 2.1, which are enforced by judicious choice of the added noise.
 Assumption 2.1.
-Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix.
-Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $k\leq q\leq p$.
-Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues.
-Let us fix:
-$\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right).$ (8)
+Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric matrix.
+Fix a target rank $k$, an intermediate rank $q\geq k$, and an iteration rank $p$, with $k\leq q\leq p$.
+Let $\bm{U}_{q}\in\mathbb{R}^{n\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues.
+Let us fix $\eta=O\left(\frac{\lambda_{q}}{\lambda_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\lambda_{k}}{\lambda_{q}}\right)},\frac{1}{\log\left(\tau n\right)}\right\}\right).$
 Assume that at every iteration $l$ of Algorithm 1, $\bm{G}_{\ell}$ satisfies, for some constant $\tau>0$:
-$\|\bm{G}_{\ell}\|_{2}$ $=O\left(\eta(\sigma_{k}-\sigma_{q+1})\right)\,,\quad\text{and}$ (9)
-$\quad\|\bm{U}_{q}^{\top}\bm{G}_{\ell}\|_{2}$ $=O\left(\eta\left(\sigma_{k}-\sigma_{q+1}\right)\frac{\sqrt{p}-\sqrt{q-1}}{\tau\sqrt{d}}\right)\,.$ (10)
-Theorem 2.4 (Private Power Method (PPM) (Balcan et al., 2016)).
-Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix.
-Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq d$.
-Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$.
+$\|\bm{G}_{\ell}\|_{2}$ $=O\left(\eta(\lambda_{k}-\lambda_{q+1})\right)\,,\quad\text{and}\quad\|\bm{U}_{q}^{\top}\bm{G}_{\ell}\|_{2}=O\left(\eta\left(\lambda_{k}-\lambda_{q+1}\right)\frac{\sqrt{p}-\sqrt{q-1}}{\tau\sqrt{n}}\right)\,.$ (4)
+We now restate the Private Power Method major result in (Balcan et al., 2016).
+Theorem 2.2 (Private Power Method (PPM), reduction to $s=1$ from the proof in Appendix C.1 from Balcan et al. (2016)).
+Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric data matrix.
+Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq n$.
+Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log(n))$.
 Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters.
-Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues.
-Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}$ gives us the following runtime-dependent convergence bound:
-$\eta=O\left(\frac{\epsilon^{-1}\max_{l}{(\|\bm{X}_{l}\|_{\max})}\sqrt{pLd\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right)$
-and also the following runtime-independent convergence bound
-$\eta=O\left(\frac{\epsilon^{-1}\|\bm{U}\|_{\max}\sqrt{pLd\log(1/\delta)\log(d)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$
-3 Analysis of Runtime-dependent Bounds
-3.1 Proposed runtime-dependent bounds
-3.1.1 Adjacency notion:
-In prior works (Hardt & Price, 2014; Balcan et al., 2016), two matrices $\bm{A}$ and $\bm{A}^{\prime}$ are considered adjacent (denoted $\bm{A}\sim\bm{A}^{\prime}$) if they differ by a change in a single element, with a Frobenius norm of at most 1:
+Let $\bm{U}_{q}\in\mathbb{R}^{n\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues.
+Then Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}$ is ($\epsilon$, $\delta$)-DP and with probability at least 0.9
+$\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}$ (5)
+[…]
+[…]
+3 Proposed Differentially Private Power Method Convergence Bounds
+In this section, we present an enhanced Differentially Private Power Method.
+First, we introduce a generalized definition of adjacency that goes beyond single-entry changes in a PSD matrix to allow for more applications.
+Then, we modify the Private Power Method to calibrate DP noise using a new, strictly tighter sensitivity bound, eliminating the $\sqrt{p}$ (associated to target rank) factor in prior work.
+We finish by showing analytically that this refinement yields strictly sharper convergence guarantees.
+Adjacency notion.
+In prior works (Hardt and Price, 2014; Balcan et al., 2016), two datasets represented by PSD matrices $\bm{A}\in\mathbb{R}^{n\times n}$ and $\bm{A}^{\prime}\in\mathbb{R}^{n\times n}$ are considered adjacent (denoted $\bm{A}\sim\bm{A}^{\prime}$) if they differ by a change in a single element, with a Frobenius norm of at most 1:
 $\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\bm{e}_{j}^{\top},$ (6)
-where $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{n}$ are canonical basis vectors of dimension $d$, $c\leq 1$ represents the magnitude of the change, and $0\leq i,j<n$.
+where $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{n}$ are canonical basis vectors, $c\leq 1$ represents the magnitude of the change, and $0\leq i,j<n$.
 This adjacency notion models a sensitive change as a modification to a single element in $\bm{A}$ to protect individual element-wise updates under Differential Privacy.
-However, since $\bm{A}$ is symmetric positive semi-definite and changes must preserve this property, this necessitates updates on the diagonal elements.
+However, since $\bm{A}$ is symmetric positive semi-definite and changes must preserve this property, this formulation restricts updates to be diagonal, hindering possible applications.
 We propose a new, more general, notion of adjacency to allow for other types of updates:
 $\bm{A}^{\prime}=\bm{A}+\bm{C},$ (7)
-where $\bm{C}\in\mathbb{R}^{n\times n}$ is a symmetric matrix representing the update, subject to the condition:
-$\sqrt{\sum_{i=1}^{n}\|\bm{C}_{i,:}\|_{1}^{2}}\leq 1.$ (13)
+where $\bm{C}\in\mathbb{R}^{n\times n}$ is a symmetric matrix representing the update, subject to $\sqrt{\sum_{i=1}^{n}\|\bm{C}_{i,:}\|_{1}^{2}}\leq 1$.
 The proposed adjacency notion is strictly more general than Equation (6).
 For example, setting $\bm{C}=c\cdot\bm{e}_{i}\bm{e}_{j}^{\top}$ recovers the original definition.
-Additionally, $\bm{C}$ can have non-zero entries on the diagonal, anti-diagonal, or any symmetric pattern, allowing for a variety of updates while maintaining symmetry.
-In the context of recommender systems where $\bm{A}=\bm{R}^{\top}\bm{R}$ and $\bm{R}$ is a user-item interaction matrix, this adjacency notion allows for element-wise changes in $\bm{R}$ (i.e.
-protect user-item interactions).
-By adopting this more general adjacency definition, we make our privacy guarantees applicable to a wider range of real-world scenarios.
-3.1.2 Improved sensitivity bound:
-The previous sensitivity bound $\Delta_{l}^{\text{prior}}$ (Hardt & Price, 2014; Balcan et al., 2016) for $\bm{A}\bm{X}^{l-1}$ was estimated by extrapolating from the case where $\bm{X}^{l-1}\in\mathbb{R}^{n\times 1}$ to the general case $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$.
+$\bm{C}$ can have non-zero entries on the diagonal, anti-diagonal, or any symmetric pattern, allowing for a variety of updates maintaining symmetry.
+In the context of recommender systems, where $\bm{A}=\bm{R}^{\top}\bm{R}$ represents the item-item similarity matrix and $\bm{R}\in\mathbb{R}^{m\times n}$ is the user-item interaction matrix, our proposed notion of adjacency enables element-wise modifications in $\bm{R}$ (i.e., protecting individual user-item interactions).
+Such changes in $\bm{R}$ propagate to multiple elements of $\bm{A}$, which could not be adequately accounted for under the previous adjacency definition (Equation (6)).
+By adopting our more general adjacency definition, we make our privacy guarantees applicable to a wider range of real-world scenarios.
+Sensitivity bound.
+The previous sensitivity bound $\Delta_{l}^{\text{prior}}$ (Hardt and Price, 2014; Balcan et al., 2016) for $\bm{A}\bm{X}^{l-1}$ defined in Equation (3) was estimated by extrapolating from the case where $\bm{X}^{l-1}\in\mathbb{R}^{n\times 1}$ to the general case $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$, leading to a dependence on $\sqrt{p}$.
 This leads to an overestimation of the sensitivity and to unnecessarily large noise addition.
-We directly analyze the change in $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$ and derive a tighter bound on the sensitivity $\Delta_{l}$.
-By using our proposed adjacency notion from Equation (12), we establish the following result (proof provided in Appendix A.1):
+We directly analyze the change in $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$ and derive a strictly tighter bound on the sensitivity $\Delta_{l}$.
+By using our adjacency notion from Equation (7), we establish the following result (proof in Appendix A.1):
 Theorem 3.1 (Improved Sensitivity Bound).
-Let $\bm{A}^{\prime}$ be defined as in Equation (12), and consider the sensitivity $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$. Then, $\Delta_{l}$ is upped-bounded as follows:
-$\Delta_{l}\leq\max_{i}\|\bm{X}_{i,:}^{l}\|_{2}\triangleq\hat{\Delta}_{l}.$ (14)
-This improved bound $\hat{\Delta}_{l}$ is generally tighter than the previous bound $\Delta_{l}^{\text{prior}}=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$, as it eliminates the dependence on $\sqrt{p}$ and directly relates to the maximum row norms of $\bm{X}^{l}$.
-By reducing the estimated sensitivity, we can add less noise while still satisfying Differential Privacy, thereby improving the utility of the algorithm.
-3.1.3 Privacy proof:
-To address these issues, we propose a result with a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2.2.
-This proof clarifies the ambiguities or errors present in the previous works (see Appendix A.2), and ensures that our privacy guarantees are met.
+Let $\bm{A}^{\prime}$ be defined as in Equation (7), and consider the sensitivity $\Delta_{l}=\sup_{\bm{A}\sim\bm{A}^{\prime}}\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$. Then,
+$\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{2}\triangleq\hat{\Delta}_{l}.$ (8)
+Privacy proof.
+Our algorithm achieves differential privacy by iteratively adding calibrated noise at each round of the power method.
+The total privacy guarantee across iterations is then derived using results from adaptive composition of DP mechanisms, as initially proposed in (Bun and Steinke, 2016), allowing us to precisely quantify the cumulative privacy loss across multiple iterative steps.
+To clarify the ambiguities or errors present in the previous works (see Appendix A.2), and ensure that our privacy guarantees are met, we propose a result with a new proof of the Differential Privacy guarantees for our overall algorithm in Theorem 3.2, whose proof is in Appendix A.2.3.
 Theorem 3.2 (Privacy proof for the PPM).
-Assuming $\delta\in(0,1)$ and $\epsilon>0$ such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$, the Private Power Method with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ satisfies ($\epsilon,\delta$)-Differential Privacy.
-3.1.4 Improved convergence bound:
-Theorem 3.3 (Bound for the noisy power method (NPM) (Balcan et al., 2016)).
-Let reals $k\leq q\leq p$.
-Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of a positive semi-definite matrix $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues and fix $\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right)$.
-If at every iteration $l$ of the NPM $\bm{G}_{\ell}$ satisfies Assumption 2.3, then after
-$L=\Theta\left(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
-iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have:
-$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta.$ (15)
-Theorem 3.4 (Improved PPM with runtime dependent bound).
-Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix.
-Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$.
-Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$.
-Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$
-Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ gives us the following runtime-dependent convergence bound:
-$\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (16)
-Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016), and can be in practice much tighter depending on $\bm{A}$.
-Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$.
-We provide the proof in Appendix A.2.2.
-4 Analysis of Runtime-independent bounds
-4.1 Proposed runtime-independent convergence bound
-We presented in Theorem 3.4 a runtime-dependent convergence bound involving $\max_{i}\|\bm{X}_{i,:}\|_{2}$, which is only known during the execution of the algorithm.
-To provide a more practical analysis, we now derive a runtime-independent convergence bound in Theorem 4.1.
-This bound is a combination of two cases:
-  • High Coherence Case: Applies to matrices with large coherence parameters, and the bound is independent of the matrix’s coherence.
-  • Low Coherence Case: Applies to matrices with low $\mu_{1}$-coherence, with the bound depending explicitly on $\mu_{1}(\bm{A})$.
-We demonstrate that our method achieves a convergence bound that is at least as tight as both of these bounds.
-We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$.
-This is likely to happen in practice, as we highlight in Section 6.1.1.
+Let $\delta\in(0,1)$ and $\epsilon>0$ such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$.
+Then, Algorithm 1 with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ is ($\epsilon,\delta$)-Differentially Private.
+Improved convergence bound.
+Building on Equation (8), we present in Theorem 3.3 a strictly tighter convergence bound than the one proposed in (Balcan et al., 2016).
+Additionally, unlike past proofs (Hardt and Price, 2014; Balcan et al., 2016), our proposed privacy proof (given in Theorem 3.2) does not restrict $\epsilon\leq 1$.
+We provide the proof in Appendix A.3.
+Theorem 3.3 (Improved PPM with Runtime-Dependent Bound).
+Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric data matrix.
+Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq n$.
+Suppose the number of iterations $L=\Theta(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log(n))$.
+Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$.
+Let $\bm{U}_{k}\in\mathbb{R}^{n\times k}$ be the top-$k$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues.
+Then Algorithm 1 is ($\epsilon$,$\delta$)-DP with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ and with probability at least 0.9
+$\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}$ (9)
+with $\text{with }\quad\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ln\log(1/\delta)\log(L)}}{\lambda_{k}-\lambda_{q+1}}\right).$ (10)
+4 Proposed Runtime-Independent Convergence Bound
+We presented in Theorem 3.3 a convergence bound involving $\max_{i}\|\bm{X}_{i:}\|_{2}$, which is only observable during the execution of the algorithm.
+To provide a more practical analysis, we now derive a runtime-independent convergence bound in Theorem 4.1 by careful bounding of $\max_{i}\|\bm{X}_{i:}\|_{2}$.
+We provide a proof in Appendix A.4.
+This bound makes it possible to have a tight analysis in two regimes:
+  • Small $\mu_{0}$-coherence, small $p$: Previously proposed in Balcan et al.
+(2016), this bound is useful in a regime with small $\mu_{0}$ when computing few eigenvectors, with a dependence on $\sqrt{p\log({n})}\cdot\mu_{0}(\bm{A})$.
+  • Larger $\mu_{0}$ or $p$: We propose a new bound tailored for the multi-dimensional power method, depending on $\mu_{1}(\bm{A})$, with a reduced dependence on the number of eigenvectors $p$.
+We note that we are likely to be in the second regime in practice, as we highlight in Section 6.1.
 Theorem 4.1.
-Improved PPM with runtime independent bound
-Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix.
-Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$.
-Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$.
-Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$ Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 gives us the following runtime-independent convergence bound:
-$\eta=O\left(\frac{\epsilon^{-1}\cdot\min(1,\mu_{1}(\bm{A}))\cdot\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (17)
+Improved PPM with Runtime-Independent Bound
+Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric data matrix.
+Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq n$.
+Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log(n))$.
+Let $\bm{U}_{q}\in\mathbb{R}^{n\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues.
+Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$ Then Algorithm 1 is ($\epsilon$,$\delta$)-DP and we have with probability at least 0.9
+$\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}$ (11)
+with $\text{with }\quad\eta=O\left(\frac{\epsilon^{-1}\cdot\min(\mu_{0}(\bm{A})\sqrt{p\cdot\log({n})},\mu_{1}(\bm{A}))\cdot\sqrt{Ln\log(1/\delta)\log(L)}}{\lambda_{k}-\lambda_{q+1}}\right).$ (12)
 5 Decentralized version
-Table 1: Overhead of using Secure Aggregation in our protocol
-Client Server
-Communication $O(log(s)+n^{2})$ $O(s\cdot log(s)+sn^{2})$
-Computation $O(log^{2}(s)+n^{2}\cdot log(s))$ $O(s\cdot log^{2}(s)+sn^{2}\cdot log(s))$
+Table 1: Per-iteration overhead: Secure Aggregation vs. Federated Learning.
+Secure Aggregation Federated Learning
+Client Server Client Server
+Communication $O\bigl{(}\log(s)+np\bigr{)}$ $O\bigl{(}s\,[\log(s)+np]\bigr{)}$ $O(np)$ $O\bigl{(}s\,np\bigr{)}$
+Computation $O\bigl{(}\log^{2}(s)+np\log(s)\bigr{)}$ $O\bigl{(}s\,[\log^{2}(s)+np\log(s)]\bigr{)}$ $O(np)$ $O\bigl{(}s\,np\bigr{)}$
 In this section, we consider a decentralized setting where the matrix $\bm{A}$ is distributed across multiple clients.
@@
 Balcan et al.
-(2016) proposed a private and federated power method using communication over public channels between clients and a server.
-However, this approach relies on Differential Privacy, since the data exchanged can be observed by everyone, which requires high levels of noise to ensure privacy.
+(2016); Guo et al.
+(2024) proposed private and federated power methods using communication over public channels between clients and a server.
+However, these approaches rely on local Differential Privacy, since the data exchanged can be observed by everyone, which requires high levels of noise to ensure privacy.
 To enhance privacy while retaining the benefits of distributed computation, we use Secure Aggregation, a lightweight Multi-Party Computation protocol.
@@
 This approach offers privacy guarantees similar to central DP and eliminates the need for a trusted curator.
-Distributed DP has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
+Distributed DP has been extensively studied in the literature (Goryczka et al., 2013; Ghazi et al., 2019; Kairouz et al., 2021; Chen et al., 2021; Wei et al., 2024).
 In distributed DP, each client adds carefully calibrated noise to their local contributions before participating in the Secure Aggregation protocol.
-The noise is designed such that the sum across clients has a variance comparable to that used in central DP, thereby achieving similar Differential Privacy guarantees without requiring a trusted aggregator, and allowing clients to keep the data locally.
-We operate under the “honest-but-curious“ threat model, where clients follow the protocol correctly but may attempt to learn additional information from received data.
-We also assume there are no corrupted or dropout users during the computation.
-For simplicity, this paper focuses on a basic implementation of distributed DP, neglecting the effects of data quantization or errors introduced by modular arithmetic.
-We refer to Kairouz et al.
-(2021) for an implementation of Distributed DP that takes these effects into account.
-Finally, it is possible to use a fully decentralized and differentially private protocol to perform averaging without relying on a central server for Secure Aggregation, as demonstrated by Sabater et al.
+The noise is designed such that the sum across clients has a variance comparable to that used in central DP, thereby achieving similar Differential Privacy guarantees without requiring a trusted aggregator, and allowing clients to keep data locally.
+We operate under the honest-but-curious threat model, where clients follow the protocol correctly but may attempt to learn information from received data.
+We also assume that there are no corrupted or dropout users during the computation.
+For simplicity, this papers neglects the effects of data quantization or errors introduced by modular arithmetic in distributed DP and refer to Kairouz et al.
+(2021) for an implementation taking this into account.
+It is also possible to use a fully decentralized DP protocol to perform secure averaging without relying on a central server, as demonstrated by Sabater et al.
 (2022).
-Adopting such approaches can further enhance decentralization while maintaining similar utility and communication costs.
+Adopting such approaches can further enhance decentralization while maintaining similar utility and communication costs for the method.
 5.1 Decentralized and private power method using distributed DP:
 We introduce a federated version of the Privacy-Preserving Power Method in Algorithm 2.
-This version reduces the noise variance by a factor of $O(sp\log d\,\|\bm{U}\|_{\infty}^{2})$ compared to the method in Balcan et al.
+This version significantly reduces the noise variance by a factor of $sp\log n\,\|\bm{U}\|_{\infty}^{2}$ compared to the method in Balcan et al.
 (2016).
 The improvement comes from the fact that we emulate centralized Differential Privacy (DP) over secure channels, avoiding the higher noise required in local DP settings with public channels, and that we use our generally tighter sensitivity bound.
-We demonstrate the equivalence between the decentralized Algorithm 2 and the centralized Algorithm 1 in Theorem 5.2 (proof provided in Appendix A.5).
-We also give a simplified definition of the Secure Aggregation algorithm.
+We now give a simplified definition of Secure Aggregation and demonstrate the equivalence between Algorithm 2 and our centralized Algorithm 1 in Theorem 5.2 (proof in Appendix A.4).
 Definition 5.1.
-We denote by $SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1{\leq}i{\leq}s\})$ the Secure Aggregation of locally held matrices $\bm{Y}_{\ell}^{(i)}$ over the set of users represented by indices $\{i|1{\leq}i{\leq}s\})$.
-It is equivalent to computing $\bm{Y}_{\ell}=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}$ over a secure channel.
-  Input: distributed matrices $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, private parameters $\epsilon,\delta$.
-  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
-  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\nu=\epsilon^{-1}\sqrt{\frac{4L\log(1/\delta)}{s}}$;
-  for $\ell=1$ to $L$ do
-      1.
+Let $SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1{\leq}i{\leq}s\})$ the Secure Aggregation of matrices $\bm{Y}_{\ell}^{(i)}$ held by users indexed by $\{i|1{\leq}i{\leq}s\})$.
+It is equivalent to computing $\bm{Y}_{\ell}=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}$ over secure channels.
+1:  Input: distributed matrices $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, private parameters $\epsilon,\delta$.
+2:  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
+3:  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\nu=\epsilon^{-1}\sqrt{\frac{4L\log(1/\delta)}{s}}$;
+4:  for $\ell=1$ to $L$ do
+5:      1.
 The central node broadcasts $\bm{X}^{\ell-1}$ to all $s$ computing nodes;2.
@@
 The central node computes QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
-  end for
+6:  end for
 Algorithm 2 Federated private power method
@@
 5.2 Communication and computation cost analysis:
-We denote by $s$ the number of shards, such that $\sum_{i=1}^{s}\bm{A}^{(i)}=\bm{A}\in\mathbb{R}^{n\times n}$.
-Table 1 presents the communication and computations cost overhead introduced by the polylogarithmic version of Secure Aggregation that we employ (Bell et al., 2020).
+Table 1 presents the per-iteration communication and computation overhead introduced by both Secure Aggregation (Bell et al., 2020) and non-secure federated aggregation.
+We can see that Secure Aggregation introduces some non-dominant logarithmic terms.
 6 Empirical comparison of the proposed bounds
-We introduced both runtime-dependent and runtime-independent convergence bounds for our algorithm.
-The runtime-dependent bound depends on the matrix $\bm{A}$ and the iteratively computed matrices $\bm{X}_{l}$, whereas the runtime-independent bound (Theorem 4.1) only depends on $\bm{A}$.
-To illustrate the practical impact of our changes, we focus on an application in recommender systems and compare how our algorithm perform in this context.
-We will then use a statistical approximation to see how the runtime-dependent bounds compare at the first step of the algorithm, regardless of the application.
+We introduced runtime-dependent and runtime-independent convergence bounds for our algorithm.
+The runtime-dependent bound depends on $\bm{A}$ and the iteratively computed $\bm{X}^{l}$, whereas the runtime-independent bound (Theorem 4.1) only depends on $\bm{A}$.
+To illustrate the practical impact of our changes, we focus on an application in recommender systems and compare how our algorithms perform in this context.
+Additionally, we use a statistical approximation to see how the bounds compare at the first step of the algorithm (sketching step), regardless of the application in Appendix C.
 6.1 Application to recommender systems
-State-of-the-art recommender systems like GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) utilize singular value decomposition (SVD) or eigenvalue decomposition as part of their algorithms.
-Specifically, they represent the dataset (comprised of user-item interactions) using a normalized adjacency matrix $\tilde{\bm{R}}$, from which they compute a normalized item-item matrix $\tilde{\bm{P}}$.
+State-of-the-art recommender systems like GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) utilize singular or eigenvalue decomposition as part of their algorithms.
+Specifically, they represent the dataset (user-item interactions) using a normalized adjacency matrix $\tilde{\bm{R}}$, from which they compute a normalized item-item matrix $\tilde{\bm{P}}$.
 To get rid of noise in $\tilde{\bm{R}}$, they compute an ideal low-pass filter based on the top-$p$ eigenvectors of $\tilde{\bm{P}}$ and apply it to $\tilde{\bm{R}}$.
-Detailed definitions are provided in Appendix A.7.
-Table 2 presents key statistics of popular recommendation datasets used in the literature.
+Detailed definitions are provided in Appendix B.
+The first four columns of Table 2 presents key statistics of popular recommendation datasets used in the literature.
 Depending on the use case, these datasets may be held by a central curator, or distributed among the users of a recommender system.
 Accordingly, either the centralized PPM (Algorithm 1) or the decentralized PPM (Algorithm 2) can be applied to compute the leading eigenvectors.
-To demonstrate the practicality of our proposed methods, we focus on a decentralized setting where each user has access to part of the dataset and collaborate with other users to compute the desired eigenvectors.
+To demonstrate the practicality of our proposed methods, we focus on a decentralized setting where each user has access to its own ratings and collaborates with other users to compute the desired eigenvectors.
+$\Delta_{l}^{prior}=\mu_{0}(\bm{A})\sqrt{p\cdot\log({n})}$ denotes the sensitivity bound proposed in (Balcan et al., 2016) while ${\hat{\Delta}_{l}}=\mu_{1}(\bm{A})$ uses our multidimensional refinement.
 Table 2: Statistics of datasets
-Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{prior}$ (existing) ${\hat{\Delta}_{l}}$ (ours)
-Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.12 ×$\times\sqrt{p}$ 1
-MovieLens 71,567 10,677 7,972,582 0.39 32.84 1.19 ×$\times\sqrt{p}$ 1
-EachMovie 74,425 1,649 2,216,887 0.49 21.89 1.33 ×$\times\sqrt{p}$ 1
-Jester 54,906 151 1,450,010 0.73 7.78 1.64 ×$\times\sqrt{p}$ 1
-6.1.1 Empirical comparison of the runtime-independent bounds
-As discussed in Section 4.1, our proposed runtime-independent bound is tighter than those proposed in Hardt & Price (2014); Balcan et al.
-(2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$.
-Table 2 shows that this condition holds for popular recommender system datasets.
+Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{prior}$ (prior) ${\hat{\Delta}_{l}}$ (ours)
+Amazon-book 52,643 91,599 2,984,108 0.33 0.99 1.12 ×$\times\sqrt{p}$ 0.99
+MovieLens 71,567 10,677 7,972,582 0.39 1.00 1.19 ×$\times\sqrt{p}$ 1.00
+EachMovie 74,425 1,649 2,216,887 0.49 1.00 1.33 ×$\times\sqrt{p}$ 1.00
+Jester 54,906 151 1,450,010 0.73 1.00 1.64 ×$\times\sqrt{p}$ 1.00
+(a) MovieLens dataset.
+(b) EachMovie dataset.
+Figure 1: Comparison of the impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}},$ where $\tilde{\bm{R}_{p}}$ represents the approximated matrix and $\bm{R}_{p}$ the original matrix.
+Results are means on 10 runs with shaded bands indicating 99% confidence intervals (computed via bootstrap), shown for the MovieLens and EachMovie datasets, with $p=32$ and $L=3$.
+Empirical comparison of the runtime-independent bounds.
+As discussed in Section 4, our proposed runtime-independent bound is tighter than those in Hardt and Price (2014); Balcan et al.
+(2016) when $\mu_{1}(\bm{A})\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({n})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$.
+The last two columns from Table 2 show that this condition holds for popular recommender system datasets.
 As we can deduce from Table 2, the proposed bound theoretically allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
-6.1.2 Practical utility of the proposed algorithms
+Practical utility of the proposed algorithms.
 We saw that the proposed runtime-independent bounds were practically tighter than the previous one for the task of interest.
-We demonstrate that our decentralized PPM can privately compute the top-$p$ eigenvectors $\bm{U}_{p}$ of the user-normalized item-item matrix $\tilde{\bm{P}}$ (proof in Appendix A.1):
+We now demonstrate that Algorithm 2 can be used to compute the top-$p$ eigenvectors $\bm{U}_{p}$ of the item-item matrix $\tilde{\bm{P}}$ under ($\epsilon$,$\delta$)-DP (proof in Appendix B.1):
 Lemma 6.1.
-Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ can compute $\bm{U}_{p}$, the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting under $(\epsilon,\delta)$-Differential Privacy.
-Figure 1: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the EachMovie dataset.
-Both GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) use $\bm{U}_{p}$ to compute the ideal low pass filter and apply it to the interaction matrix $\bm{R}$, giving a filtered interaction matrix $\bm{R}_{p}$.
-To illustrate the practicality of our proposed method, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using either our proposed decentralized PPM, or the PPM versions proposed by  Balcan et al.
+Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ can approximate $\bm{U}_{p}$, the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting under $(\epsilon,\delta)$-Differential Privacy.
+Both GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) use $\bm{U}_{p}$ to compute the ideal low pass filter and apply it to the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$.
+To illustrate the practicality of our proposed method, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using either our decentralized PPM, the previous PPM versions in  Balcan et al.
+(2016) or FedPower from Guo et al.
 (2024).
-We then use $\tilde{\bm{U}}_{p}$ to compute an approximation of $\bm{R}_{p}$ (noted $\tilde{\bm{R}_{p}}$), and compute the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to each version.
-Figure 1 illustrates the impact of the privacy parameter $\epsilon$ on the approximation error for the EachMovie dataset, with $p=32$ and $L=3$.
-Our decentralized PPM achieves a relatively low approximation error (approximately 1/10 of the random guessing error) for practical values of $\epsilon$ (5-10).
-In contrast, the methods from Balcan et al.
-(2016) require $\epsilon$ values of at least $10^{3}$ to achieve comparable errors, which does not provide meaningful privacy protection.
-Additional comparisons and details are provided in Appendix A.7.1.
-6.2 Matrix-agnostic comparison of the runtime-dependent bounds
-We analyze the tightness of our proposed runtime-dependent bound compared to the previous bound in a matrix-agnostic manner, focusing on the first iteration of the algorithm.
-At iteration $l=0$, the leading singular vectors are initialized as $\bm{X}^{0}=\text{Q}(\bm{\Omega})$, where $\bm{\Omega}\sim\mathcal{N}(0,1)^{n\times k}$ and $\text{Q}(\bm{\Omega})$ denotes the orthonormal matrix obtained from the QR decomposition of $\bm{\Omega}$.
-Since $\bm{X}^{0}$ is independent of the matrix $\bm{A}$, it allows us to assess the relative tightness of the proposed bound $\hat{\Delta}_{0}$ compared to the previous bound $\Delta_{0}^{prior}$ at the first step.
-Let $r(k,n)$ be the ratio of the expected values of the two bounds, as a function of the desired number of eigenvectors $k$ and the matrix dimension $n$:
-$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{prior}]}{\mathbb{E}[\hat{\Delta}_{0}]}.$ (18)
-6.2.1 Theoretical approximation of $r(k,n)$:
-We derive an asymptotic approximation of $r(k,n)$ using Approximation 6.2 (detailed in Appendix A.4).
-Approximation 6.2.
-[…]
-$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
-$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
-6.2.2 Empirical approximation of $r(k,n)$:
-Figure 2: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=8000$ and $k$ ranging between 64 and 4000, with a step of 64.
-Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm.
-The theoretical ratios are estimated using Theorem 6.2.
-To compare the tightness of our proposed bound $\hat{\Delta}_{0}$ to the previous bound $\Delta_{0}^{prior}$, we perform both theoretical and empirical analyses at the first iteration, where the noise scaling depends only on the random initialization $\bm{X}^{0}$.
-Since $\bm{X}^{0}$ is independent of $\bm{A}$, we can estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$ by sampling random Gaussian matrices $\bm{X}^{0}$.
-Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$.
-We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$.
-We use Theorem 6.2 to approximate it theoretically.
-To measure it empirically, we compute $\Delta_{0}^{prior}$ and $\hat{\Delta}_{0}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm.
-We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$.
-We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples.
-We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 6.2.
-Figure 2 and Figure 3 (in Appendix) presents the comparison between empirical (blue dots) and theoretical (green line) estimates of $r(k,n)$ for $n=8000$ and $n=12000$.
-The empirical ratios are calculated from the averages over the $t=5$ runs for each value of $k$.
-The results indicate that both empirically and theoretically the proposed noise scaling $\hat{\Delta}_{l}$ is much tighter than $\Delta_{l}^{prior}$ at the first step of the algorithm.
-Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice.
-The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
-7 Conclusion and limitations
-We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems.
-We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees.
-By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, maintaining the efficiency and privacy of the centralized version but adapting it for distributed environments.
-However, our privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic.
-In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
-8 Impact Statement (required by ICML)
-Our contributions directly address the need for privacy-preserving algorithms in the context of large-scale machine learning systems, where sensitive personal or organizational data is often at stake.
-The decentralized implementation ensures that the method can be applied in some real-world environments where data is distributed across multiple users or devices, making it particularly suitable for recommender systems and other distributed applications.
-This work is expected to have implications for privacy-conscious machine learning in real-world systems, to perform large-scale data analysis while maintaining robust privacy protections.
+We then use $\tilde{\bm{U}}_{p}$ to compute an approximation of $\bm{R}_{p}$ (denoted by $\tilde{\bm{R}_{p}}$), and compute the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to each version.
+We provide experimental details in Appendix B.2.
+Figure 1 illustrates the impact of the privacy parameter $\epsilon$ on the approximation error for the EachMovie and MovieLens datasets, with $p=32$ and $L=3$, clearly showing the advantage of our method over prior work.
+Our decentralized PPM achieves relative approximation error of $\approx 1/10$ for values of $\epsilon\in(5,10)$ for EachMovie and $\epsilon\approx 20$ for MovieLens.
+In contrast, other decentralized methods (Balcan et al., 2016; Guo et al., 2024) require $\epsilon$ to be of the order of at least $10^{3}$ to achieve comparable errors, which does not seem to provide meaningful privacy protection.
+[^3] We refer to Dwork et al.
+(2019) for more practical details on how to set $\epsilon$ and provide additional comparisons and details in Appendix B.1.
+(Guo et al., 2024) yields the worst relative approximation errors.
+We hypothesize that this is because it uses worst case, non-adaptive sensitivity bounds for DP, as opposed to our proposed method and those of Balcan et al.
+(2016).
+For a fixed approximation error, both our centralized and decentralized methods yield $\epsilon$ roughly four times smaller than required by the centralized method of Balcan et al.
+(2016), demonstrating that our propositions significantly strengthen privacy guarantees.
+7 Conclusion
+We presented differentially private versions of the centralized and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems.
+We introduced a new sensitivity bound, which we show theoretically and empirically to improve the accuracy of the method while ensuring privacy guarantees.
+By employing Secure Aggregation in a decentralized setting, we can reduce the noise introduced for privacy, maintaining the efficiency and privacy of the centralized version but adapting it for distributed environments.
+Our methods could enable organizations (from healthcare networks to social platforms) to extract useful structure from data without compromising individual records.
+Limitations.
+Our privacy guarantees require no dropout from participants to ensure distributed DP, which may not always be realistic.
+It would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially converge faster.
 Appendix A Proofs of Results
@@
 Proof.
-We note $\bm{A}^{\prime}$ a matrix adjacent to $\bm{A}$ using (12).
+We denote by $\bm{A}^{\prime}$ a matrix adjacent to $\bm{A}$ using (7):
 $\bm{A}^{\prime}=\bm{A}+\bm{C},$ (13)
@@
 $\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\hat{\Delta}_{l}.\\ \end{split}$ (16)
-Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\hat{\Delta}_{l}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
+Although the proposed update model (7) is more general than when $\bm{A}^{\prime}$ is defined using (6), the proposed bound $\hat{\Delta}_{l}$ is also generally tighter than the bound proposed in Hardt and Price (2014); Balcan et al.
+(2016).
 A.2 Privacy proof
 A.2.1 Note on related privacy proofs:
-Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by Hardt & Price (2014), Balcan et al.
-(2016), Hardt & Roth (2013), Hardt & Roth (2012).
+Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by Hardt and Roth (2012, 2013); Hardt and Price (2014); Balcan et al.
+(2016).
 These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP.
 Specifically, Balcan et al.
-(2016) references the privacy proof from Hardt & Price (2014), which builds upon the privacy proof from Hardt & Roth (2013).
-The proof therein is also closely related to the one of Hardt & Roth (2012).
-However, both Hardt & Roth (2013) (Theorem 2.4) and Hardt & Roth (2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped.
+(2016) references the privacy proof from Hardt and Price (2014), which builds upon the privacy proof from Hardt and Roth (2013).
+The proof therein is also closely related to the one of Hardt and Roth (2012).
+However, both Hardt and Roth (2013) (Theorem 2.4) and Hardt and Roth (2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped.
 This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all.
@@
 (2010).
-Moreover, Lemma 3.4 of Hardt & Roth (2013) misuses Theorem 2.4.
+Moreover, Lemma 3.4 of Hardt and Roth (2013) misuses Theorem 2.4.
 Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration.
 By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
-A.2.2 Proof of Theorem 3.2
+A.2.2 Zero-Concentrated Differential Privacy (zCDP):
+A randomized algorithm $\mathcal{M}$ is said to satisfy $\rho$-zero-Concentrated Differential Privacy (zCDP) if for all neighboring datasets $D_{1}$ and $D_{2}$, and for all $\alpha\in(1,\infty)$, the following holds:
+$D_{\alpha}(\mathcal{M}(D_{1})\|\mathcal{M}(D_{2}))\leq\rho\alpha\,,$ (17)
+where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees).
+The following lemma, introduced by Bun and Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
+Lemma A.1 (Gaussian Mechanism (Proposition 1.6 (Bun and Steinke, 2016))).
+Let $f:X^{n}\to\mathbb{R}$ be a sensitivity-$\Delta$ function.
+Consider the mechanism (randomized algorithm) $M:X^{n}\to\mathbb{R}$, defined as $M(x)=f(x)+Z_{x}$ where for each input $x$, $Z_{x}$ is independently drawn from $\mathcal{N}(0,\sigma^{2})$.
+Then $M$ satisfies $\left(\frac{\Delta^{2}}{2\sigma^{2}}\right)$-zCDP.
+The next lemma, which is a generalized version of a result presented by Bun and Steinke (2016), explains how a randomized algorithm, constructed by recursively composing a sequence of zCDP-satisying randomized algorithms, also satisfies zCDP.
+Lemma A.2 (Adaptive composition (Generalization from Lemma 2.3 of (Bun and Steinke, 2016))).
+Let $M_{1}:X^{n}\to Y_{1}$, $M_{2}:X^{n}\times Y_{1}\to Y_{2}$, …, $M_{L}:X^{n}\times Y_{1}\times\dots\times Y_{L-1}\to Y_{L}$ be randomized algorithms.
+Suppose $M_{i}$ satisfies $\rho_{i}$-zCDP as a function of its first argument for each $i=1,2,\dots,L$.
+Let $M^{\prime\prime}:X^{n}\to Y_{L}$, constructed recursively by:
+$M^{\prime\prime}(x)=M_{L}(x,M_{L-1}(x,\dots,M_{2}(x,M_{1}(x))\dots)).$ (18)
+Then $M^{\prime\prime}$ satisfies $(\sum_{i=1}^{L}\rho_{i})$-zCDP.
+A.2.3 Proof of Theorem 3.2
 Proof.
 We can see lines 4-5 of Algorithm 1 as a sequential composition ($M$) of $L$ Gaussian Mechanisms.
-By Lemma 2.1, each mechanism $M_{i}$ satisfies ($\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP where $\Delta_{i}$ is the $\ell_{2}$-sensitivity of the function associated to mechanism $\Delta_{i}$ and $\sigma_{i}^{2}$ is the variance of the noise added with the Gaussian Mechanism.
-By Lemma 2.2, the composition of mechanisms $M=(M_{1},...,M_{i},...,M_{L})$ satisfies ($\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP.
+By Lemma A.1, each mechanism $M_{i}$ satisfies ($\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP where $\Delta_{i}$ is the $\ell_{2}$-sensitivity of the function associated to mechanism $\Delta_{i}$ and $\sigma_{i}^{2}$ is the variance of the noise added with the Gaussian Mechanism.
+By Lemma A.2, the composition of mechanisms $M=(M_{1},...,M_{i},...,M_{L})$ satisfies ($\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP.
 Let $\rho\triangleq\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$.
@@
 $=\frac{\epsilon^{2}}{{8\log(1/\delta)}}.$ (23)
-By Proposition 1.3 of Bun & Steinke (2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$.
+By Proposition 1.3 of Bun and Steinke (2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$.
 Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
@@
 $\iff\delta\leq\exp{(-\frac{\epsilon}{4})},$ (31)
-which is a reasonable assumption, as in practice $\epsilon=O(1)$ and $\delta\ll\frac{1}{d}$ where $d$ is the number of records to protect.
+which is a reasonable assumption, as in practice $\epsilon=O(1)$ and $\delta\ll\frac{1}{d}$, where $d$ is the number of records to protect.
 In our case, $d=n^{2}$ because we run the privacy-preserving power method on $\bm{A}\in\mathbb{R}^{n\times n}$.
@@
 A.3 Proof of Theorem 3.3
+The following theorem from Balcan et al. (2016) will be useful in our proof:
+Theorem A.3 (Bound for the noisy power method (NPM) (Balcan et al., 2016)).
+Let $k\leq q\leq p$ be positive integers.
+Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of a positive semi-definite matrix $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{d}\geq 0$ denote its eigenvalues and fix $\eta=O\left(\frac{\lambda_{q}}{\lambda_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\lambda_{k}}{\lambda_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right)$.
+If at every iteration $l$ of the NPM $\bm{G}_{\ell}$ satisfies Assumption 2.1, then after
+$L=\Theta\left(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
+iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have:
+$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}.$ (32)
+We now provide the proof of Theorem 3.3:
 Proof.
-According to (Hardt & Price, 2014), if $\bm{G}_{l}\sim\mathcal{N}(0,\sigma_{l})^{d\times p}$, then with probability p=99/100 we have:
+According to Hardt and Price (2014), if $\bm{G}_{l}\sim\mathcal{N}(0,\sigma_{l})^{d\times p}$, then with probability 99/100 we have:
 $\begin{split}&\max_{l=1}^{L}\|\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{d\cdot\log(L)}\,,\\ &\max_{l=1}^{L}\|\bm{U}^{\top}\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{p\cdot\log(L)}\,.\end{split}$ (33)
-We can therefore satisfy the noise conditions of Theorem 3.3 with probability $p$ if we choose $\eta=\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}$, giving us:
-$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}\,.$ (41)
-Which leads us to the following Theorem by injecting $\sigma_{l}=\hat{\Delta}_{l}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
-A.4 Derivation of Approximation 6.2
-We provide here the derivation for Approximation 6.2:
-Proof.
-Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries.
-Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{i}sQ$ orthogonal and $\bm{R}$ upper triangular).
-By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$.
-Then by Theorem 2.2.1 of Chikuse (2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries.
-We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty.
-For a matrix $\bm{A}$, we note as $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation.
-Then we have:
-$\bm{Q}_{ij}^{2}$ $=(\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})_{ij}^{2}$ (42)
-$=(\bm{Z}^{\circ 2}((\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})^{\circ 2})_{ij}$ (43)
-$=(\bm{Z}^{\circ 2}(\bm{Z}^{\top}\bm{Z})^{-1})_{ij}.$ (44)
-We remark that each element of $\bm{Z}^{\circ 2}$ is distributed as a chi-squared variable with one degree of freedom, therefore $n\cdot\bm{Q}_{ij}^{2}$ is distributed as an i.i.d $F(1,n)$ random variable by definition.
-For large $n$, we can approximate these $F$ variables as Gaussians with mean $\mu=\frac{n}{n-2}$ and variances $\sigma^{2}=\frac{2n^{2}(n-1)}{(n-2)^{2}(n-4)}$.
-Therefore $\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{\mu}{n},\frac{\sigma^{2}}{n^{2}})$ and:
-$k\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k^{2}\sigma^{2}}{n^{2}})$ $\sum_{j}\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k\sigma^{2}}{n^{2}}).$ (45)
-We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 from Massart (1896).
-We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$.
-We assume that the variances of $\Delta_{l}^{prior}$ and $\hat{\Delta}_{l}$ (as $n$ and $k$ are large) are small and therefore:
-$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\mathbb{E}[(\Delta_{l}^{prior})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
-$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\mathbb{E}[(\hat{\Delta}_{l})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
-Figure 3: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=12000$ and $k$ ranging between 64 and 4000, with a step of 64.
-Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm.
-The theoretical ratios are estimated using Theorem 6.2.
-A.5 Proof of Theorem 4.1
+We can therefore satisfy the noise conditions of Theorem A.3 with probability 99/100 if we choose $\eta=\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\lambda_{k}-\lambda_{q+1}}$, giving us:
+$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\lambda_{k}-\lambda_{q+1}}\,,$ (34)
+which leads us to the statement of the theorem by injecting $\sigma_{l}=\hat{\Delta}_{l}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$. ∎
+A.4 Proof of Theorem 4.1
 We provide here the proof for Theorem 4.1: See 4.1
 Proof.
-First bound: We know by Section 2.4 from Woodruff et al.
-(2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1.
-Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
-Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM.
-As proposed in Hardt & Price (2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar.
-Therefore we have:
-$<x_{c},x_{o}>$ $=<\sum_{i=1}^{n}\alpha_{c}^{i}u_{i},\sum_{j=1}^{n}\alpha_{o}^{i}u_{j}>$ (48)
-$=\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}<u_{i},u_{j}>$ (By bilinearity) (49)
-$=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}.$ (Orth. columns of $\bm{U}$) (50)
-$\bm{X}$ has orthonormal columns, so if $c=o$, then $<x_{c},x_{o}>=\sum_{i=1}^{n}{(\alpha_{c}^{i})}^{2}=1$.
-Otherwise, we have $<x_{c},x_{o}>=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}=0$.
-We can therefore define a matrix $\bm{B}$ with orthonormal columns such that $\bm{X}=\bm{U}\bm{B}$ and $\bm{B}_{jc}=\alpha_{c}^{j}$.
-We can therefore write any row $i$ of $\bm{X}$ as $\bm{X}_{i:}=\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}$.
-Therefore
-$\|\bm{X}_{i:}\|_{2}$ $=\|\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (51)
-$\leq\sum_{j=1}^{p}\|\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (52)
-$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|\|\bm{B}_{j:}\|_{2}$ (53)
-$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|.$ (Orth. columns of $\bm{B}$) (54)
-We then have $\max_{i}\|\bm{X}_{i:}\|_{2}\leq\max_{i}(\sum_{j=1}^{p}|\bm{U}_{ij}|)=\|\bm{U}\|_{\infty}=\mu_{1}(\bm{A})$.
-We proposed two bounds for $\|\bm{X}_{i:}\|_{2}$, which are useful in two different cases. We can however note that:
-$\|\bm{X}_{i:}\|_{2}\leq 1\text{ and }\|\bm{X}_{i:}\|_{2}\leq\mu_{1}(\bm{A})$ (55)
-$\Rightarrow\|\bm{X}_{i:}\|_{2}\leq\min(1,\mu_{1}(\bm{A})),$ (56)
-which leads us to Theorem 4.1 giving us a runtime-independent bound.
-A.6 Proof of Theorem 5.2
+We showed in Equation that $\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}$.
+This quantity depends on values computed during the execution of the algorithm.
+We now show that we can bound $\max_{i}\|\bm{X}_{i:}^{l}\|_{F}$ with a runtime-independent bound.
+Without loss of generality and to simplify notation, we use $\bm{X}$ to denote any matrix $\bm{X}_{l}$ computed during the execution of the Private Power Method.
+Let $\bm{X}_{:c}=x_{c}$ denote a column of $\bm{X}$.
+As $\bm{A}$ is an Hermitian matrix, by the spectral theorem, we have $\bm{A}=\bm{U}\bm{D}\bm{U}^{\top}$, where $\bm{U}$ is unitary (with orthonormal columns) and $\bm{D}$ is diagonal.
+As the columns of $\bm{U}$ form a complete basis for $\mathbb{R}^{n}$, we can write any column $x_{c}$ of $\bm{X}$ as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$, where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar.
+We can then write:
+$\langle x_{c},x_{e}\rangle$ $=\langle\sum_{i=1}^{n}\alpha_{c}^{i}u_{i},\sum_{j=1}^{n}\alpha_{e}^{j}u_{j}\rangle$ (35)
+$=\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{c}^{i}\alpha_{e}^{j}\langle u_{i},u_{j}\rangle$ (36)
+$=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{e}^{i}.$ (Orth. columns of $\bm{U}$) (37)
+$\bm{X}$ is the “$Q$“ matrix constructed from a Gram-Schmidt QR decomposition, it has therefore orthonormal columns by definition.
+Therefore, if $c=e$, then $\langle x_{c},x_{e}\rangle=\sum_{i=1}^{n}{(\alpha_{c}^{i})}^{2}=1$.
+Otherwise, we have $\langle x_{c},x_{e}\rangle=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{e}^{i}=0$.
+The key is then to notice that we can define a matrix $\bm{B}\in\mathbb{R}^{n\times p}$ with orthonormal columns such that $\bm{X}=\bm{U}\bm{B}$ and $\bm{B}_{jc}=\alpha_{c}^{j}$.
+We recall that any matrices $\bm{H}$ and $\bm{J}$, we have $\|\bm{H}\bm{J}\|_{F}\leq\|\bm{H}\|_{2}\|\bm{J}\|_{F}$.
+Then, we can bound the norm of any row of $\bm{X}_{i:}$ as:
+$\|\bm{X}_{i:}\|_{F}$ $=\|\bm{X}_{i:}^{\top}\|_{F}\,,$ (38)
+$=\|\bm{B}^{\top}\bm{U}_{i:}^{\top}\|_{F}\,,$ (39)
+$\leq\|\bm{B}^{\top}\|_{2}\|\bm{U}_{i:}^{\top}\|_{F}.$ (40)
+$\bm{B}$ has orthonormal columns by construction, therefore
+$\|\bm{B}\|_{2}$ $=\sqrt{\|\bm{B}^{\top}\bm{B}\|_{2}}$ (41)
+$=\sqrt{\|\bm{I}\|_{2}}$ (42)
+$=1.$ (43)
+We then have:
+$\|\bm{X}_{i:}\|_{F}$ $\leq\|\bm{B}^{\top}\|_{2}\|\bm{U}_{i:}^{\top}\|_{F}$ (44)
+$\leq\|\bm{U}_{i:}^{\top}\|_{F}$ (45)
+$\leq\|\bm{U}_{i:}\|_{F}.$ (46)
+Additionally as $\bm{X}_{i:}\in\mathbb{R}^{1\times n}$, $\max_{i}\|\bm{X}_{i:}\|_{2}=\max_{i}\|\bm{X}_{i:}\|_{F}\leq\max_{i}(\|\bm{U}_{i:}\|_{F})=\mu_{1}(\bm{A})$.
+Note: By Section 2.4 from Woodruff (2014), for any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$, $\|v\|_{2}\leq 1$.
+As $\bm{U}$ has orthonormal columns by construction, $\max_{i}\|\bm{U}_{i:}\|_{2}\leq 1$.
+We can therefore bound $\max_{i}\|\bm{X}_{i:}\|_{2}$ as:
+$\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}\|_{2}\leq\mu_{1}(\bm{A})\leq 1.$ (47)
+Injecting this bound in Theorem 3.3 leads us to Theorem 4.1, giving us a runtime-independent bound. ∎
+A.5 Proof of Theorem 5.2
 Proof.
@@
 where $\bm{G}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\cdot(s\nu^{2}))^{n\times p}$, and we have $s\nu^{2}=\sigma^{2}$ by definition, completing the equivalence proof. ∎
-A.7 Application to recommender systems:
+Appendix B Application to recommender systems:
 Let $s$ be the number of users in our system and $n$ the number of items.
@@
 $\begin{split}\bm{P}^{\prime}&=\tilde{\bm{R}}^{\top}\tilde{\bm{R}}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}}).\\ \end{split}$ (49)
-To simplify our analysis, we consider $\bm{I}$ public and do not use item-wise normalization in the computation of ideal low pass filter, leaving it for future work.
+To simplify our analysis, we consider $\bm{I}$ public and do not use item-wise normalization in the computation of the ideal low pass filter, leaving it for future work.
 We therefore define $\tilde{\bm{R}}=\bm{U}^{-\frac{1}{2}}\bm{R}$ and $\tilde{\bm{P}}=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})$.
-Lemma A.1.
+Lemma B.1.
 We can use Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ to compute the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting with under a $(\epsilon,\delta)$-Differential Privacy guarantee.
 Proof.
-We note $\tilde{\bm{P}}_{ij}$ as the element of $\tilde{\bm{P}}$ at line $i$ and column $j$ and $d_{user}(u)=\sum_{i=0}^{n-1}r_{ui}$. We can write $\tilde{\bm{P}}_{ij}$ as such:
-$\begin{split}\tilde{\bm{P}}_{ij}&=((\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}))_{ij}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{T})_{i,*}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=((\bm{U}^{-\frac{1}{2}}\bm{R})_{*,i})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=\sum_{u=0}^{s-1}\frac{1}{\sqrt{d_{user}(u)}}r_{ui}\frac{1}{\sqrt{d_{user}(u)}}r_{uj}\\ &=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}.\\ \end{split}$ (59)
-By noticing that $(\bm{R}^{\top}\bm{R})_{ij}=\sum_{u}r_{ui}\cdot r_{uj}$, then we can deduce that $\tilde{\bm{P}}=\sum_{u}\frac{1}{d_{user}(u)}\bm{R}_{u}^{\top}\bm{R}_{u}$.
+Let $\tilde{\bm{P}}_{ij}$ denote the element of $\tilde{\bm{P}}$ at row $i$ and column $j$ and let $d_{user}(u)=\sum_{i=0}^{n-1}r_{ui}$. We can write $\tilde{\bm{P}}_{ij}$ as:
+$\begin{split}\tilde{\bm{P}}_{ij}&=((\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}))_{ij}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top})_{i,*}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=((\bm{U}^{-\frac{1}{2}}\bm{R})_{*,i})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=\sum_{u=0}^{s-1}\frac{1}{\sqrt{d_{user}(u)}}r_{ui}\frac{1}{\sqrt{d_{user}(u)}}r_{uj}\\ &=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}.\\ \end{split}$ (50)
+By noticing that $(\bm{R}^{\top}\bm{R})_{ij}=\sum_{u}r_{ui}\cdot r_{uj}$, we can deduce that $\tilde{\bm{P}}=\sum_{u}\frac{1}{d_{user}(u)}\bm{R}_{u}^{\top}\bm{R}_{u}$.
 Therefore $\tilde{\bm{P}}$ is partitioned among $s$ users as described in Section 5.
@@
 $\bm{A}^{\prime}_{ij}=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)-1}\cdot r_{ui}^{\prime}\cdot r_{uj}^{\prime},$ (52)
-where $r_{ui}^{\prime}=r_{ui}$ except for one user-item interaction, i.e.
-$r_{vk}=1$ but $r_{vk}^{\prime}=0$.
+where $r_{ui}^{\prime}=r_{ui}$ except for one user-item interaction, i.e., $r_{vk}=1$ but $r_{vk}^{\prime}=0$.
 Let $\bm{C}=\bm{A}-\bm{A}^{\prime}$.
@@
 $=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (56)
-$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}.$ (66)
+$\quad\quad+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}.$ (57)
 We have:
@@
 $\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}$ $=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (67)
-$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (80)
+$\quad\quad+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (68)
 $=\frac{2(d_{user}(v)-1)}{d_{user}(v)^{2}}+1$ (69)
@@
 $\implies\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{2}.$ (73)
-A.7.1 Approximation errors comparisons:
-As explained in Section 6.1.2, GF-CF and BSPM (with its parameter $T_{b}=1$) use $\bm{U}_{p}$ to compute the ideal low pass filter and filter the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$.
+B.1 Approximation errors comparisons:
+As explained in Section 6.1, GF-CF and BSPM (with its parameter $T_{b}=1$) use $\bm{U}_{p}$ to compute the ideal low pass filter and filter the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$.
 Indeed, we have:
 $\bm{R}_{p}=\bm{R}\cdot\bm{I}^{-\frac{1}{2}}\bm{U}_{p}\bm{U}_{p}^{\top}\bm{I}^{\frac{1}{2}}.$ (74)
-To study the impact of Differential Privacy on our system, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using our proposed decentralized PPM, or the PPM versions of Balcan et al.
-(2016).We have:
+To study the impact of Differential Privacy on our system, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using our proposed decentralized PPM, the PPM versions of Balcan et al.
+(2016) or FedPower from Guo et al.
+(2024).
+We have:
 $\tilde{\bm{R}}_{p}=\bm{R}\cdot\bm{I}^{-\frac{1}{2}}\tilde{\bm{U}}_{p}\tilde{\bm{U}}_{p}^{\top}\bm{I}^{\frac{1}{2}}.$ (75)
@@
 (2023).
-We use $p=32$ to have acceptable approximation error (on the order of 0.02-0.05 for numbers bounded between 0 and 1) for reasonable values of $\epsilon$ (5-10).
-We showed the impact of the Differential Privacy parameter $\epsilon$ on the approximation errors for the EachMovie dataset in Figure 1.
-We get similar results for the MovieLens and Jester datasets in Figures 5 and  4.
-Figure 4: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the Jester dataset.
-Figure 5: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the MovieLens dataset.
+We use $p=32$ to have acceptable approximation error for reasonable values of $\epsilon$ (5-10).
+We use the synchronous version of FedPower (Guo et al., 2024) to simplify the comparison, i.e., we set their parameter $\mathcal{I}_{T}=L$.
+We note that FedPower could be improved by also using Secure Aggregation and therefore reducing the noise necessary for DP.
+It might also benefit from our sensitivity analysis in the synchronous setting.
+We showed the impact of the Differential Privacy parameter $\epsilon$ on the approximation errors for the MovieLens and EachMovie datasets in Figures 1(a) and 1(b).The trends for the approximation errors on the Jester dataset are quite similar, as shown in Figure 2.
+We however note that all methods perform better on this dataset.
+We hypothesize that this is because the Jester dataset is much more dense compared to the EachMovie and MovieLens datasets, hence the ratio of magnitude of the noise added due to DP compared to the magnitude of the elements of the item-item matrix is smaller on this dataset.
+Figure 2: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed on 10 runs with 99% confidence intervals (computed via bootstrap), with $p=32$ and $L=3$ on the Jester dataset.
+B.2 Experimental Details
+Datasets.
+We benchmark our method on four standard recommendation datasets: MovieLens-10M, EachMovie, Jester, and Amazon-book.
+For Amazon-book, we use the publicly available 80%/20% train/test splits; for EachMovie, MovieLens-10M and Jester we use the same train/test proportions and therefore apply a 80%/20% split at the user level (each user retains at least one test interaction).
+Table 3 summarizes the post-processing statistics.
+Table 3: Dataset statistics after preprocessing: number of users, items, interactions.
+Dataset Users Items Interactions
+MovieLens-10M 71,567 10,677 7,972,582
+EachMovie 74,425 1,649 2,216,887
+Jester 54,906 151 1,450,010
+Amazon-book 52,643 91,599 2,984,108
+Hyperparameters.
+We fix the number of power iterations to $L=3$ to correspond to PyTorch’s svd_lowrank approximate basis subroutine (Algorithm 4.3 in (Halko et al., 2011)).
+No further hyperparameter tuning is performed.
+Experimental Protocol and Runs.
+Each “run” uses a different Pseudo Random Number Generator seed for the Gaussian initialization $\bm{G}_{0}$.
+We perform 10 runs per setting (for one given $\epsilon$ value, all methods are initialized with the same Gaussian matrix), varying only the PRNG seed across runs.
+Confidence Intervals.
+We report the mean and a 99% confidence interval over the 10 runs, estimated via nonparametric bootstrap (1 000 resamples).
+Computational Resources.
+All experiments ran on a 16-core CPU. Figure 1 (10 runs × 14 $\epsilon$ values × 5 methods) requires $\approx 30$ min of wall-clock time.
+Appendix C Matrix-agnostic comparison of the runtime-dependent bounds
+We analyze the tightness of our proposed runtime-dependent bound compared to the previous bound in a matrix-agnostic manner, focusing on the first iteration of the algorithm.
+At iteration $l=0$, the leading singular vectors are initialized as $\bm{X}^{0}=\text{Q}(\bm{\Omega})$, where $\bm{\Omega}\sim\mathcal{N}(0,1)^{n\times k}$ and $\text{Q}(\bm{\Omega})$ denotes the orthonormal matrix obtained from the QR decomposition of $\bm{\Omega}$.
+Since $\bm{X}^{0}$ is independent of the matrix $\bm{A}$, it allows us to assess the relative tightness of the proposed bound $\hat{\Delta}_{0}$ compared to the previous bound $\Delta_{0}^{prior}$ at the first step.
+We note that when the randomized power method runs for only one step, it is akin to sketching (Halko et al., 2011).
+Let $r(k,n)$ be the ratio of the expected values of the two bounds, as a function of the desired number of eigenvectors $k$ and the matrix dimension $n$, i.e.
+$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{prior}]}{\mathbb{E}[\hat{\Delta}_{0}]}.$
+Theoretical approximation of $r(k,n)$. We derive an asymptotic approximation of $r(k,n)$ using Approximation C.1.
+Approximation C.1.
+[…]
+$\mathbb{E}[\Delta_{l}^{prior}]^{2}$ $\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)\text{ and }\mathbb{E}[\hat{\Delta}_{l}]^{2}$ $\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).\vspace{-5mm}$ (76)
+We provide here the derivation for Approximation C.1:
+Proof.
+Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries.
+Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{Q}$ is orthogonal and $\bm{R}$ upper triangular).
+By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$.
+Then by Theorem 2.2.1 of Chikuse (2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries.
+We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty.
+For a matrix $\bm{A}$, we denote by $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation.
+Then we have:
+$\bm{Q}_{ij}^{2}$ $=(\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})_{ij}^{2}$ (77)
+$=(\bm{Z}^{\circ 2}((\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})^{\circ 2})_{ij}$ (78)
+$=(\bm{Z}^{\circ 2}(\bm{Z}^{\top}\bm{Z})^{-1})_{ij}.$ (79)
+We remark that each element of $\bm{Z}^{\circ 2}$ is distributed as a chi-squared variable with one degree of freedom, and therefore $n\cdot\bm{Q}_{ij}^{2}$ is distributed as an i.i.d $F(1,n)$ random variable by definition.
+For large $n$, we can approximate these $F$ variables as Gaussians with mean $\mu=\frac{n}{n-2}$ and variances $\sigma^{2}=\frac{2n^{2}(n-1)}{(n-2)^{2}(n-4)}$.
+Therefore $\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{\mu}{n},\frac{\sigma^{2}}{n^{2}})$ and:
+$k\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k^{2}\sigma^{2}}{n^{2}})\,,$ $\sum_{j}\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k\sigma^{2}}{n^{2}}).$ (80)
+We can approximate the expectation of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 from Massart (2007).
+We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$.
+We assume that the variances of $\Delta_{l}^{prior}$ and $\hat{\Delta}_{l}$ are small (because $n$ and $k$ are large) and therefore:
+$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\mathbb{E}[(\Delta_{l}^{prior})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}\,,$ (81)
+$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\mathbb{E}[(\hat{\Delta}_{l})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (82)
+Empirical approximation of $r(k,n)$.
+To compare the tightness of our proposed bound $\hat{\Delta}_{0}$ to the previous bound $\Delta_{0}^{prior}$, we perform both theoretical and empirical analyses at the first iteration, where the noise scaling depends only on the random initialization $\bm{X}^{0}$.
+Since $\bm{X}^{0}$ is independent of $\bm{A}$, we can estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$ by sampling a random Gaussian matrix $\bm{X}^{0}$.
+Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$.
+We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$.
+We use Approximation C.1 to approximate it theoretically.
+To measure it empirically, we compute $\Delta_{0}^{prior}$ and $\hat{\Delta}_{0}$ based on $\bm{X}^{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm.
+We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$.
+We claim that $t=5$ is enough to see a general trend, as we have estimates for many ($k,n$) couples.
+We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation C.1.
+(a) $n=8000$.
+(b) $n=12000$.
+Figure 3: Comparison of empirically (green dots) and theoretically (blue line) estimated $r(k,n)$ ratios for $k$ ranging between 64 and 4000 with a step of 64.
+Empirical ratios are estimated using $t=5$ runs of the first step of our algorithm, while theoretical ratios are based on Theorem C.1.
+Results are shown for two different values of $n$: 8000 and 12000.
+Figures 3(a) and 3(b) present the comparison between empirical (blue dots) and theoretical (green line) estimates of $r(k,n)$ for $n=8000$ and $n=12000$, respectively.
+The empirical ratios are calculated from the averages over the $t=5$ runs for each value of $k$.
+The results indicate that both empirically and theoretically the proposed noise scaling $\hat{\Delta}_{l}$ is much tighter than $\Delta_{l}^{prior}$ at the first step of the algorithm.
+Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice.
+The proposed bound is tighter by a multiplicative factor on the first step and therefore drastically reduce the impact of the noise introduced by DP at the first iteration.
+By noting that the power method is usually run for very few steps ($L$ is usually in the range of $1$-$5$), this result complements our general convergence bounds on the overall algorithm derived in Theorems 3.3 and 4.1 and gives further intuition on the tightness and usefulness of our proposed sensitivity bound.